### v1

# MnLargeSymbols’164 MnLargeSymbols’171

Abstract
Recent advancements in moiré engineering motivate study of the behavior of strongly-correlated electrons subject to substantial orbital magnetic flux. We investigate the triangular lattice Hofstadter-Hubbard model at one-quarter flux quantum per plaquette and a density of one electron per site, where geometric frustration has been argued to stabilize a chiral spin liquid phase intermediate between the weak-coupling integer quantum Hall and strong-coupling 120 antiferromagnetic phases. In this work, we use Density Matrix Renormalization Group methods and analytical arguments to analyze the compactification of the Hofstadter-Hubbard model to cylinders of finite radius. We introduce a glide particle-hole symmetry operation which for odd-circumference cylinders, we show, is spontaneously broken at the quantum Hall to spin liquid transition. We further demonstrate that the transition is associated with a diverging correlation length of a charge-neutral operator. For even-circumference cylinders the transition is associated with a dramatic quantitative enhancement in the correlation length upon threading external magnetic flux. Altogether, we argue that the 2+1D CSL-IQH transition is in fact continuous and features critical correlations of the charge density and other spin rotationally-invariant observables.
Introduction—The Kalmeyer-Laughlin chiral spin liquid (CSL) is one of the earliest and most important examples of topological order [1, 2, 3] and has been extensively studied in spin systems, both numerically [4] and through partonic mean-field approaches [5, 6, 7]. Recently, considerable efforts have been devoted to identifying electronic settings in which the CSL may emerge to connect more directly to solid state experiments, in which electrons—rather than spins—are the fundamental degrees of freedom and may be far from their Mott limit. These systems enable the exploration of the Mott transition of strongly-interacting electrons in strong magnetic fields, a regime that is becoming experimentally accessible both at non-zero [8, 9] and zero Zeeman coupling [10, 11]. They also set the stage for the study of charge fluctuation-driven transitions out of the CSL phase into other phases such as superconductors [12], exotic charge-density waves [13], or quantum Hall states [14, 11].
Figure 1: (a) Upper panel: The Hofstadter-Hubbard model on the 2D triangular lattice, with magnetic translations $T_{x,y}$, on-site Hubbard $U$ interaction, and flux per triangular plaquette $\Phi_{\triangle}$ indicated. Lower left panel: “Hofstadter butterfly” representation of spectrum of non-interacting model in the plane of single-particle energy $E$ and flux per plaquette $\Phi_{\triangle}$. $\Phi_{\triangle}=\pi/2$ is the unique flux for which the Hamiltonian enjoys a doubled “magnetic” unit cell and a particle-hole symmetry $\mathcal{P}$ relating opposite $C=\pm 1$ bands [lower right panel]. (b) A finite segment of the infinite-length, circumference-$L_{y}$ cylinder geometry, threaded by $\Phi_{\text{ext}}$ external magnetic flux. With the chosen “YC” cylindrical geometry, the translation $T_{y}$ is orthogonal to the cylinder axis and $x\in\mathbb{Z}$ indexes rings. (c) Expected 2+1D phase diagram at half-filling. As interactions increase, the spin-degenerate integer quantum Hall (IQH) ground state undergoes a quantum phase transition into a Kalmeyer-Laughlin chiral spin liquid (CSL). By studying both even and odd finite-circumference cylinders, we shed light on the existence and nature of this transition.
A model electronic system thought to host both conventional and CSL electronic phases is the triangular Hofstadter-Hubbard model (see Fig. 1a) in the regime of strong orbital magnetic flux, where the CSL may be favored energetically by both geometric frustration [15, 16] and the explicit breaking of time-reversal symmetry [17, 7]. The Hofstadter-Hubbard model is parametrized by an on-site interaction $U$ and a nearest neighbor hopping with Peierls phases corresponding to a flux of $\Phi_{\triangle}$ per elementary plaquette. When $\Phi_{\triangle}=\pi/2$ per triangle, this system realizes a spin-unpolarized integer quantum Hall (IQH) insulator at small values of the Hubbard interaction, with $\sigma_{xy}=2e^{2}/h$, obtained by filling the lower topological sub-band in both spin sectors. This state possesses a gap to spin and charge excitations, is topological and has a spin-degenerate charged edge mode. At very large $U$ and density of one electron per site, the state is a Mott insulator: the leading term in the $t/U$ expansion around the infinite-$U$ Mott limit shows that the low-energy physics is governed by the conventional antiferromagnetic nearest-neighbor $S=1/2$ Heisenberg model with $J_{H}\sim t^{2}/U$, which has $120^{\circ}$ antiferromagnetic order at $T=0$, so that the ground state is non-topological, with fully-gapped charge excitations and soft (Goldstone) spin excitations.
Carrying out the $t/U$ expansion to higher order suggests the existence of a CSL phase intermediate between the Mott/AF and IQH phases: the next correction is a third-order chiral term $S\cdot S\times S$ with coefficient $J_{\triangle}\sim t^{3}\sin\Phi_{\triangle}/U^{2}$. In the presence of the magnetic flux this term explicitly breaks both time-reversal and parity symmetry [17, 7]. Numerical studies of the $J_{H}$-$J_{\triangle}$ spin Hamiltonian have indeed revealed a CSL ground state for large enough $J_{\triangle}/J_{H}$ [18, 19], although at these values the accuracy of truncating the $t/U$ expansion at third order may be questioned. Very recently Kuhlenkamp et al. presented cylinder iDMRG evidence of an intermediate CSL phase for $U\gtrsim 10t$ at fluxes $\Phi_{\triangle}=\pi/3,\pi/2$ [11]. The CSL phase was identified by the absence of charge pumping as flux was threaded through the cylinder and the presence of appropriate degeneracies in the entanglement spectrum.
Supposing one accepts the existence of an intermediate CSL phase, the question of the nature of the IQH-CSL transition arises. A parton theory in which one decomposes the electron into a bosonic chargon $b$ carrying the charge and a fermionic spinon $f$ carrying the spin index, $c_{\sigma}=bf_{\sigma}$ has been proposed [14, 12]. This decomposition requires an emergent gauge field $a$, corresponding to the relative phase between the partons, under which they carry opposite unit $U(1)$ charge. In both the IQH and CSL phases the spinon occupies both $C=1$ magnetic sub-bands descending from the electronic bandstructure, while at the electronic IQH-CSL phase transition the chargon and the gauge field together undergo a superconductor-Mott phase transition [12]. The dynamical gauge field $a$ is responsible not only for producing the correct topological response on both sides of the transition, but also for gapping out the charge sector in the IQH phase through the Anderson-Higgs mechanism [20, 21], and allowing for spinon deconfinement in the CSL [22, 23, 24]. The parton representation and related slave-particle numerical studies on honeycomb and square lattices  [14, 25] have provided a physically intuitive, yet incomplete mean-field picture of the putative continuous transition. The iDMRG investigation of Ref. [11] revealed the IQH-CSL transition only as a crossover indicated by a weak maximum in a correlation length. Further investigation of the CSL phase and the IQH-CSL transition is therefore needed.
In this work we use a compactification to an infinite-length, finite-radius cylinder [see Fig. 1(b)] in combination with a symmetry analysis with large-scale iDMRG simulations [26, 27] on cylinders of circumference $L_{y}=3$ to $6$ to characterize the nature of the IQH and CSL states, as well as the transition between them. Our symmetry analysis is based on an operator that is a combination of particle-hole interchange and translation, which we show explicitly is spontaneously broken for interactions above a critical interaction strength for odd-circumference cylinders. We further show that the broken symmetry occurs together with the degeneracy pattern in the entanglement spectrum expected for a CSL phase. We present numerical evidence that this transition is in fact continuous and argue that in the limit of infinite cylinder radius, this transition evolves into a continuous 2+1D IQH-CSL quantum phase transition featuring critical fluctuations in the charge density and other spin rotationally-invariant observables.
Figure 2: CSL symmetry-breaking at odd circumference. (a) Graphical illustration of the action of the particle-hole symmetry operator $\mathcal{P}$ which transforms the gauge flux through a ring of an odd-$L_{y}$ cylinder, from $\phi$ to $\pi-\phi$. (b) Upper panel: In an odd-$L_{y}$ CSL ground state, the alternating pattern $+\pi/2,-\pi/2,\dots$ of these gauge fluxes spontaneously breaks $\mathcal{P}T_{x}$ symmetry, whose action is equivalent to the threading of $\pi$ gauge flux. Lower panel: the low-lying entanglement eigenvalues at cuts between two pairs of rings, which not only exhibit the approximate degeneracy expected for the CSL at each $S_{z}$ [see circled numerals], but also transform as half-integer spin representations only at every other cut, signalling SSB of $\mathcal{P}T_{x}$. (c) In contrast, the internal gauge flux configuration [upper panel] and entanglement spectrum [lower panel] of the IQH state are consistent with it being a unique, symmetry-preserving ground state.
Model, symmetry operator and cylinder compactification—We study the Hofstadter-Hubbard model on the triangular lattice, defined pictorially in Fig. 1(a). In this model, electrons hop between nearest-neighboring sites and experience an Aharonov-Bohm phase of $\Phi_{\triangle}$ upon orbiting a triangular plaquette. Note that this model focuses on the consequences of the orbital motion, and sets the Zeeman coupling to zero. We fix $\Phi_{\triangle}=\pi/2$, which endows the Hamiltonian with a unitary particle-hole (PH) symmetry $\mathcal{P}$ acting as [28]
$\mathcal{P}c^{\dagger}_{x,y}\mathcal{P}^{-1}=(-1)^{x+y}(i\sigma_{y})c_{x,y},\quad\mathcal{P}i\mathcal{P}^{-1}=+i,$ (1)
on fermionic operators, where we have attached a conventional spin rotation $\sigma_{y}$ to the naive definition of particle-hole transformation in order for $\mathcal{P}$ to leave the spin operator $\bm{S}=\frac{1}{2}c^{\dagger}\bm{\sigma}c$ invariant (see SM for details [29]). The flux induces a magnetic translation algebra $T_{x}T_{y}=-T_{y}T_{x}$ that we represent in a gauge where
$T_{x}c^{\dagger}_{x,y}T_{x}^{-1}=(-1)^{y}c^{\dagger}_{x+1,y},\quad T_{y}c^{\dagger}_{x,y}T_{y}^{-1}=c^{\dagger}_{x,y+1}.$ (2)
The commuting translations $(T_{x})^{2},T_{y}$ define a larger “magnetic” unit cell, whose resulting two sub-bands are related by $\mathcal{P}$ and carry opposite Chern number $C=\pm 1$. We fix the filling to $n=1$ electron per lattice site, which in the absence of interactions corresponds to full filling of the lowest spin-degenerate magnetic sub-band.
We now present the dimensional reduction of the problem to a cylinder. The key results of our study rely on the different symmetry breaking patterns that occur for even and odd circumferences $L_{y}$, which stem from generic properties of magnetic fields (both external and internal) in the chosen geometry. To illustrate how this arises, let us first consider the closed “tin can”-shaped surface specified by a pair of adjacent rings at $x$ and $x+1$. This surface is the union of $2L_{y}$ triangles on the outside of the cylinder, as well as an $L_{y}$-sided polygonal surface at $x$ and another at $x+1$ whose edges are the rings. Since $\pi/2$ external flux penetrates each outer triangle, Gauss’s law requires that the external magnetic flux through the two adjacent rings differ by $\pi L_{y}$ [30, 31]. These ring fluxes are therefore uniform when $L_{y}$ is even, but are staggered when $L_{y}$ is odd, which explicitly breaks $T_{x}$ symmetry.
Let us momentarily focus on the odd case and fix the sequence of ring fluxes to alternate between 00 and $\pi$. In this case, the system nonetheless respects a combined translation-like symmetry, namely the product of $\mathcal{P}$ and $T_{x}$—which we refer to as PH-glide—in spite of each being individually broken by the compactification. The IQH state, being adiabatically connected to the (unique) ground state of the non-interacting Hamiltonian, is invariant under PH-glide symmetry.
We argue, however, that the CSL phase must break $\mathcal{P}T_{x}$ spontaneously when $L_{y}$ is odd. To see this, note that the CSL is a Mott insulating phase. Thus, we may consider the limit where charge fluctuations are frozen out, e.g. by increasing $U\to\infty$ while holding the three-spin chiral term at a fixed favorable strength. Since the particle-hole operation is transparent to the spin operator, $\mathcal{P}\bm{S}\mathcal{P}^{-1}=\bm{S}$, this results in a spin model with a pure translation symmetry $t_{x}$.00footnotetext: We note a similar observation in a simpler 1+1D model [32]. Since each ring contains an odd number of spin-$1/2$ moments, the Lieb-Schulz-Mattis theorem requires that any gapped ground state be at least two-fold degenerate [33], e.g., by spontaneously breaking $t_{x}$. One way to understand the translation symmetry breaking of the CSL, in particular, is by leveraging its anyonic content. Namely, the properties of the CSL under translations are captured by viewing it has hosting a spinon in each unit cell of the triangular lattice [34, 35, 36]. Because these spinons have semionic statistics, the action of $t_{x}$ translation—which moves $L_{y}$ many spinons across any given cut—necessarily permutes the two minimally-entangled CSL ground states [37] when $L_{y}$ is odd [38, 39]. In other words, the only $t_{x}$-symmetric ground states are long-range-entangled CSL “cat” states, implying that translation is spontaneously broken [40].
Figure 3: Evidence for the IQH-CSL phase transition. (a) Upper panel: Sketch of three rings of cylinder at indicated $x$ values, with solid and dashed lines showing along which staggered-Heisenberg order parameter $\mathcal{O}_{\delta\mathrm{Heis}}$ [defined in Eq. (3)] changes under glide particle-hole symmetry $\mathcal{P}T_{x}$ on odd-circumference cylinders. Lower panel: Magnitude of the staggered Heisenberg order parameter plotted against $U/t$ for cylinder widths indicated and with $\mathcal{P}T_{x}$ preserving and broken phases indicated by the check and cross symbols. Consistent with the recovery of this symmetry at the putative 2+1D critical point, the magnitude of the order parameter [bottom left] decreases substantially from $L_{y}=3$ to $5$. (b-d) Correlation length in the $(Q,S_{z},k_{y})=\bm{0}$ charge sector, plotted against interaction strength, for a range of MPS bond dimensions [increasing light to dark] and extrapolated to infinite bond dimension [gray points]; the insets specify the external flux threading each of the cylinder rings. (b) At circumference $L_{y}=5$, the $\mathcal{P}T_{x}$ symmetry-breaking is associated with a diverging correlation length. (c) While both the IQH and CSL states are fully-symmetric on even-$L_{y}$ cylinders, the dominant correlation length is consistently charge-neutral, peaking strongly for $\Phi_{\text{ext}}=\pi/2$ external flux threading the cylinder (see inset). The maximum extrapolated value of the correlation length increases from $L_{y}=4$ to $6$, consistent with an approach to a continuous transition in the thermodynamic limit.
It is illustrative to phrase the above discussion in terms of operations on the internal gauge field introduced in the parton construction. For simplicity, let us focus on saddle point configurations of the gauge field on each side of the transition, ignoring fluctuations. In both the CSL and IQH ground states, the gauge flux through the triangular surface plaquettes should match the external flux, $\phi_{\triangle}=\pi/2$ [6, 7, 12]. Thus, the only remaining gauge-invariant degree of freedom is the gauge flux through the cylinder rings; assuming $T_{x}^{2}$ is unbroken and $L_{y}$ is odd, we may parameterize the two fluxes as $(\phi,\phi+\pi)$, differing by $\pi$ due to Gauss’s law, as invoked above. Since the minimally-entangled CSL ground states are adiabatically connected to the Mott limit in which $\mathcal{P}$ acts trivially, their ring flux patterns are restricted to be $(\pi/2,-\pi/2)$ and $(-\pi/2,\pi/2)$, which are insensitive to the action of $\mathcal{P}$ on ring fluxes, namely $\phi\to\pi-\phi$ (Fig. 2a). We readily see that these two configurations are related by $\mathcal{P}T_{x}$, exposing the two-fold degeneracy and SSB of the CSL (Fig. 2b, top) [41]. In the IQH phase, on the other hand, the Meissner effect pins together the internal and external gauge fields [42]. The resulting flux pattern $(0,\pi)$ maps into itself under $\mathcal{P}T_{x}$, consistent with a unique, gapped ground state (Fig. 2c, top).
While the small $U$ IQH phase does not break the glide-PH symmetry and the intermediate $U$ CSL phase has a two-fold symmetry breaking pattern on odd circumference cylinders, the large $U$ $120^{\circ}$ phase breaks translational symmetry (including the glide PH) in a threefold spatial pattern.
Numerical observation of phase transition—We first comment on evidence for the phases themselves. The primary diagnostic of chiral topological behavior is the behavior of the charge-resolved entanglement spectrum (ES) [43, 44] at a bipartition of the system into two semi-infinite halves, in this case between any pair of adjacent rings [40, 11]. We exploit the relationship between the physical and entanglement edges [45, 43, 46], namely that the low-lying states in the entanglement spectrum correspond to those of the edge CFT [47, 48, 49, 50].
Within the bulk-boundary correspondence, the CSL is expected to host gapless edges described at low energies by a free bosonic conformal field theory [51]. The counting of low-energy states as a function of momentum in each conserved charge sector can be obtained combinatorially: for fixed $n\in\mathbb{Z}_{\geq 0}$ and associated momentum $q_{y}=2\pi n/L_{y}$, one considers a linearly dispersing mode and imagines exciting states from below to above the Fermi level. Note that these bosonic objects are charge and spin-neutral but carry momentum. The counting sequence is namely the number of integer partitions of $n$: $(1,1,2,3,5,\dots).$ On the other hand, the IQH phase has two independent bosonic edge modes. Here, the total momentum must be the sum of those of the two flavors. Working out the “two-color” integer partitions of $n$ yields the counting $(1,2,5,10,20\dots).$ These expectations are borne out in our numerics; see Fig. 2(b,c).
We now investigate the transition. Consistent with the expectations above, we confirm that the CSL phase spontaneously breaks $\mathcal{P}T_{x}$ symmetry on odd-$L_{y}$ cylinders. We first find that the bipartite ES in the CSL phase exhibits a two-ring periodicity, which implies SSB. In contrast, the ES of the IQH is fully invariant under the action of $\mathcal{P}T_{x}$, as shown in Fig. 2(c). Second, by scanning $U/t$ and inspecting the transfer matrix spectrum, we identify a diverging correlation length for the accessible odd-circumference cylinder geometries, namely $L_{y}=3,5$. The divergence is found to belong to the $(Q,S_{z},k_{y})=\bm{0}$ charge sector (see Fig. 3b for the $L_{y}=5$ and the SM [29] for $L_{y}=3$). The spontaneous breaking of glide-PH symmetry is the source of this quasi-(1+1)D criticality. To substantiate this, we identified an order parameter $\mathcal{O}_{\delta\mathrm{Heis}}=\sum_{m}\mathcal{O}^{\delta\mathrm{Heis}}_{2m}$, where
$\mathcal{O}^{\delta\text{Heis}}_{x}=L_{y}^{-1}\sum_{y}\bm{S}_{x,y}\cdot\bm{S}_{x+1,y}-\bm{S}_{x+1,y}\cdot\bm{S}_{x+2,y}.$ (3)
The operator $\mathcal{O}_{\delta\mathrm{Heis}}$ is both charged under $\mathcal{P}T_{x}$ and condenses in the CSL phase [see Fig. 3(a)], signaling SSB. We remark that threading external flux $\Phi_{\text{ext}}=\pi/2$ through an odd-circumference cylinder reduces the transition to a crossover, owing to the fact that flux threading explicitly breaks glide-PH symmetry unless $\Phi_{\text{ext}}\in\pi\mathbb{Z}$ (see SM [29]).
On the even-circumference geometries $L_{y}=4$ and $6$, we reproduce the weak crossover, with a correlation length of $\sim 2$ lattice sites, found by Kuhlenkamp et al. for $L_{y}=6$ and $\Phi_{\text{ext}}=0$ [11]. This is consistent with the absence of translation symmetry-breaking in the even-circumference Mott limit. Upon threading $\Phi_{\text{ext}}=\pi/2$ external flux through each of these systems, we observe dramatically larger values of the correlation length for $U$ near $U_{c}$ (Fig. 3), notably in the same neutral sector in which the odd-circumference correlation length diverges [see Fig. 3(c,d)]. We also perform an extrapolation to infinite bond dimension following Ref. [52]; see SM for details [29]. The maximal extrapolated value is $\xi=6.6$ at $L_{y}=4$, increasing substantially to $\xi=9.1$ at $L_{y}=6$. Consistent with the view that accurate estimation of the 2+1D gap and other properties requires consideration of all threaded cylinder fluxes [53], we conjecture that the $\Phi_{\text{ext}}=\pi/2$ threaded system is more reflective of the thermodynamic limit than the unthreaded system, and that larger even-$L_{y}$ geometries will exhibit an increasingly-sharp sequence of crossovers toward 2+1D critical behavior in the charge-neutral sector.
Figure 4: Density-density [left panel] and spin-spin [right panel] connected correlations for odd-circumference cylinders with $U$ tuned to the respective critical points, $U_{*}(L_{y}=3)=11.1$ and $U_{*}(L_{y}=5)=11.6$. Both $L_{y}=5$ correlation functions are scaled down by 100 for ease of viewing. (a) Density-density correlations decay nearly algebraically at the IQH-CSL transition, pointing to the presence of gapless charge-neutral excitations in the 2+1D limit. (b) In contrast, correlations of the spin density decay exponentially with no apparent enhancement from $L_{y}=3$ to $5$, suggesting that the excitations are spin-rotationally-invariant singlets.
Remarkably, the critical fluctuations at the odd-$L_{y}$ critical points manifest not only in the order parameter, but also in the charge density. This is seen manifestly in the connected density-density correlations $\langle n(x)n(0)\rangle$, with $n(x)$ averaged over the ring $x$, which we find to decay approximately algebraically, i.e., linearly on a log-log scale [see Fig. 4(a)] However, estimation of the 2+1D critical exponents from these data is hampered by the fact that the critical features of the thermodynamic limit need only show up in the regime $1\ll x\ll L_{y}$, which is inaccessible for the presently-tractable system sizes. While PH-glide is no longer broken for $L_{y}=4,6$ and the operator Eq. (3) no longer condenses (see Fig. 3a), we also observe that their enhanced correlation length is accompanied by an apparent algebraic plateau in $\langle n(x)n(0)\rangle$ over the scale of the correlation length (see SM [29]).
Curiously, we do not observe critical behavior in the odd-$L_{y}$ connected correlations of the spin density $S_{z}$ [Fig. 4(b)], which is also neutral under the conserved charge and spin $U(1)$ symmetries. However, since $S_{z}$ may be rotated into $S_{x}$ by a global spin rotation, under which the CSL is invariant, this is consistent with our not observing a diverging correlation length in the $(Q,S_{z})=(0,\hbar)$ sector of the transfer matrix. Altogether, under the assumption that this behavior persists for the entire sequence of odd circumferences, our observations suggest that the 2+1D critical point hosts gapless, charge-neutral spin-singlet excitations.
We remark that these data are consistent with the partonic view that the transition into the IQH corresponds to chargon condensation. While the chargon is not gauge invariant and may not develop conventional long-range connected correlations, at the critical point the combination of the chargon, anti-chargon and gauge string is a gapless neutral excitation [42]. Interestingly, the power law correlations presented in Fig. 4 show that the physical charge density operator, which has nonzero overlap with the gauge-invariant chargon density $b^{\dagger}(x)b(x)$, produces a state with a nonzero overlap with the critical excitations when applied to the ground state, raising the intriguing possibility that the critical fluctuations can be experimentally observed.
Discussion—By combining numerical data on both even and odd-circumference cylinder geometries, we have provided evidence that the CSL-IQH transition is in fact continuous in the Hofstadter-Hubbard model at $\pi/2$ flux per triangle and that critical properties are already manifest at the numerically-accessible circumferences $3\leq L_{y}\leq 6$. Our finding that critical fluctuations appear in the charge density correlation function, which may be probed by experimental or numerical measurement of the structure factor, raise the possibility of experimental and numerical investigation of this transition.
This paradigmatic Hofstadter-Hubbard system provides a uniquely minimal and experimentally-relevant setting [10, 11] in which to study a fundamental phase transition at the interface of spin liquid and quantum Hall physics. As such, a more complete understanding of this transition is an important task for future work. Quantitatively, our investigation clarifies the possible observable, numerical signatures that may be used to probe the phase transition in a variety of numerical methods, such as AFQMC, VMC, or fermionic PEPS, which may be able to access larger system sizes or spectral signatures of spin-charge separation [54, 55, 56].
An important broader message of this work is that the ability to extrapolate 2+1D physics from dimensionally-reduced numerical data—namely on infinite cylinder geometries—is greatly enhanced by carefully considering the manner in which candidate ground states and their phase transitions respond to dimensional reduction and to external fields that become boundary effects in the thermodynamic limit. Applying such techniques more systematically, it would be valuable to characterize the contexts in which UV features of Mott phases, namely spin liquids, may be leveraged to better characterize their charge fluctuation-induced phase transitions.
Acknowledgements—We acknowledge useful discussions with E. Altman, S. Chatterjee, O. Gauthé, M. Knap, C. Kuhlenkamp, J.Y. Lee, C. Liu, F. Pollmann, Z. Weinstein, and especially X.Y. Song. The Flatiron Institute is a division of the Simons Foundation. AJM was supported in part by Programmable Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0019443. S.D. acknowledges support from the NSERC PGSD fellowship. M.Z. and S.D. were supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (QSA). This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8683 to T.S. Calculations were performed using the TeNPy Library [57].
Supplementary Material: “Chiral Spin Liquid and Quantum Phase Transition in the Triangular Lattice Hofstadter-Hubbard Model”
Appendix A Hamiltonian construction at all rational fluxes
For the underlying triangular lattice — with the “YC” orientation on the cylinder (where the $\hat{y}$ site-to-site distance is the NN distance $a$) — we take the following Bravais vectors:
[…]
whose corresponding reciprocal lattice vectors are
[…]
SI Choosing the magnetic unit cell
The following discussion applies to both the YC and XC lattice geometries. Let the flux per parallelogram be
$\Phi=\frac{p}{q}\Phi_{0},$ (3)
where $\Phi_{0}=\frac{2\pi\hbar}{e}\sim 2\pi$ in units where $\hbar/e=1.$ The above definition of flux is convenient for defining the geometry, but we briefly remark that the flux $\Phi/2$ per triangle — which is the smallest traversable plaquette — is the quantity with respect to which physical behavior (e.g., the Hofstadter spectrum) must be $\Phi_{0}$-periodic. In any case, the corresponding magnetic field strength is
$B=\frac{\Phi}{A_{\text{par}}}=\frac{p\Phi_{0}/q}{\frac{\sqrt{3}}{2}a^{2}}=\frac{2p\Phi_{0}}{\sqrt{3}q}a^{-2}.$ (4)
Regardless of the parity of $q$, there will be $q$ magnetic sublattice sites in the magnetic unit cell, which we label by $\tau=0,\dots,q-1.$ When $q$ is even, the magnetic Bravais vectors are
$\bm{R}_{1}=q\bm{a}_{1}-\frac{q}{2}\bm{a}_{2}\propto\hat{x},\qquad\bm{R}_{2}=\bm{a}_{2}\propto\hat{y},\qquad(q\ \text{even}),$ (5)
whose total area $A_{\text{muc}}$ (for either YC or XC) is precisely that of $2q$ triangles of area $\frac{\sqrt{3}}{4}a^{2}.$ The associated reciprocal lattice vectors are
$\bm{G}_{1}=\frac{1}{q}\bm{b}_{1}\propto\hat{x},\qquad\bm{G}_{2}=\frac{1}{2}\bm{b}_{1}+\bm{b}_{2}\propto\hat{y},\qquad(q\ \text{even}).$ (6)
As a sanity check, the total of the magnetic Brillouin zone (mBZ) is then
$A_{\text{mBZ}}=|\bm{G}_{1}\times\bm{G}_{2}|=(2\pi)^{2}\frac{2}{q\sqrt{3}a^{2}}=\frac{(2\pi)^{2}}{A_{\text{muc}}}.$ (7)
When $q$ is odd, the enclosing magnetic unit cell again contains exactly $2q$ triangles, but we take
$\bm{R}_{1}=q\bm{a}_{1}-\frac{q-1}{2}\bm{a}_{2},\qquad\bm{R}_{2}=\bm{a}_{2}\propto\hat{y},,\qquad(q\ \text{odd}),$ (8)
whose total area is once again
$|\bm{R}_{1}\times\bm{R}_{2}|=q|\bm{a}_{1}\times\bm{a}_{2}|=\frac{q\sqrt{3}}{2}a^{2}=A_{\text{muc}}.$ (9)
The reciprocal lattice vectors are now
$\bm{G}_{1}=\frac{1}{q}\bm{b}_{1},\qquad\bm{G}_{2}=\frac{q-1}{2q}\bm{b}_{1}+\bm{b}_{2},\qquad(q\ \text{odd}),$ (10)
which also satisfy the inverse area relationship, Eq. (7).
In Fig. LABEL:fig:YC_c2zsymm_locations, we illustrate the choice of magnetic unit cell for both odd and even $q$ for the YC triangular lattice boundary conditions. For both even and odd denominator $q,$ note that the enclosing magnetic unit cell contains exactly $2q$ triangles, so $2\pi p\in 2\pi\mathbb{Z}$ total flux. In addition, we position the sublattice sites $\tau$ within the magnetic unit cell in such a way that all the sites within magnetic unit cell $\bm{R}$ transform under $C_{2z}$ into sites in $-\bm{R}$. In particular, the sites in the home magnetic unit cell transform into one another in a clean way. We achieved this by further partitioning the odd-$q$ geometries into two subclasses. In Fig. LABEL:fig:XC_c2zsymm_locations, we make the same specifications for the XC boundary conditions.
SII Hamiltonian and choice of vector potential
Placing the underlying triangular lattice on the cylinder, let the resulting system have circumference $L_{y}$, i.e., the points $\bm{r}\sim\bm{r}+L_{y}\bm{a}_{2}$ are geometrically identified. Let the lattice sites be indexed by
$\bm{r}=x\bm{a}_{1}+y\bm{a}_{2},$ (11)
which defines our coordinate system, where $(x,y)\in\mathbb{Z}^{2}.$ Since $\bm{a}_{2}$ is purely vertical in both geometries, then $x$ indexes the rings, and $y$ is a integer spatial index within each ring. To generate complex hoppings and modify the translation algebra to a magnetic translation algebra, we introduce an external $U(1)$ vector potential. We assume that it is linear and that its curl is the magnetic field:
$\bm{A}(\bm{r})=M\bm{r},\qquad\nabla\times\bm{A}(\bm{r})=B\hat{z},$ (12)
where $B$ is as in Eq. (4). The hoppings are obtained from the vector potential via the Peierl’s substitution, for which we adopt the following sign convention:
[…]
where the second equality is due to linearity of $\bm{A}(\bm{r})$. Our convention is that the single-particle Hamiltonian is (note the minus sign):
$\hat{h}=-\sum_{\bm{r}^{\prime},\bm{r}}c^{\dagger}(\bm{r}^{\prime})t_{\bm{r}^{\prime}\leftarrow\bm{r}}c(\bm{r}),$ (14)
where we only include nearest-neighbour hopping on the triangular lattice, for which
$t_{\bm{r}^{\prime}\leftarrow\bm{r}}=te^{i\bm{A}\left(\frac{\bm{r}^{\prime}+\bm{r}}{2}\right)\cdot(\bm{r}^{\prime}-\bm{r})}.$ (15)
To constrain the form Eq. (12), we remark that the terms $x\hat{x}$ and $y\hat{y}$ are curl-free, whereas $Bx\hat{y}$ and $-By\hat{x}$ — the standard Landau gauges on the rectangular lattice — have the correct curl. We therefore reason that the most general permissible $M$ is given by
$M=B\begin{pmatrix}\alpha&\gamma-1\\ \gamma&\beta\end{pmatrix},$ (16)
where $\alpha,\beta,\gamma\in\mathbb{R}$. We will now demand that the choice of vector potential is compatible with our chosen geometry of magnetic Bravais vectors. Specifically, we ask that the Peierl’s hoppings generated by $\bm{A}$ are periodic in the chosen Bravais vectors. This allows us to place the system easily on a torus or cylinder, provided the system tiles an integer number of magnetic unit cells. Quantitatively, we demand that the RHS of Eq. (13) is invariant under shifts of the argument of $\bm{A}$ by the magnetic Bravais vectors $\bm{R}_{1}$ and $\bm{R}_{2}$, for all bond vectors $\bm{r}^{\prime}-\bm{r}$. By linearity of the vector potential, and since the bond vectors are linear combinations of $\bm{a}_{1}$ and $\bm{a}_{2}$, it suffices to require that
$\bm{a}_{i}\cdot\bm{A}(\bm{R}_{j})\in 2\pi\mathbb{Z},$ (17)
for all $i,j$. Following Ref. [58], an elegant solution can be found in terms of the reciprocal vectors of the underlying triangular lattice. We take the following since $\bm{b}_{1}\propto\hat{x}$ for both YC and XC geometries, and we want the vector potential to be independent of $y$:
$\bm{A}(\bm{r})=\frac{\Phi}{(2\pi)^{2}}(\bm{r}\cdot\bm{b}_{1})\bm{b}_{2}.$ (18)
where $\Phi$ is the flux per parallelogram. Noting the vector identity
[×$\left[\nabla\times\left((\bm{r}\cdot\bm{Q})\bm{P}\right)\right]_{i}$ $=\epsilon_{ijk}\partial_{j}\left(\bm{r}_{\ell}\bm{Q}_{\ell}\bm{P}_{k}\right)=\epsilon_{ijk}\bm{Q}_{j}\bm{P}_{k}=(\bm{Q}\times\bm{P})_{i},$ (19)
then
$\bm{B}(\bm{r})$ =×$=\nabla\times\bm{A}(\bm{r})=\frac{\Phi}{(2\pi)^{2}}\bm{b}_{1}\times\bm{b}_{2}=\frac{\Phi}{(2\pi)^{2}}\frac{(2\pi)^{2}}{A_{\text{par}}}\hat{z}=\frac{\Phi}{A_{\text{par}}}\hat{z}.$ (20)
Moreover, it can be shown to have the required periodicity property:
$\bm{a}_{i}\cdot\bm{A}(\bm{R}_{j})$ $=\frac{\Phi}{(2\pi)^{2}}(\bm{R}_{j}\cdot\bm{b}_{1})\bm{b}_{2}\cdot\bm{a}_{i}=\delta_{i,2}\frac{2\pi p/q}{2\pi}(\bm{R}_{j}\cdot\bm{b}_{1})=\delta_{i,2}\frac{p}{q}\bm{R}_{j}\cdot\bm{b}_{1},$ (21)
which clearly vanishes when $j=2.$ Otherwise,
$\bm{a}_{i}\cdot\bm{A}(\bm{R}_{1})$ $=\delta_{i,2}\frac{p}{q}\left(\begin{cases}\left(q\bm{a}_{1}-\frac{q}{2}\bm{a}_{2}\right)\cdot\bm{b}_{1}&q\text{ even}\\ \left(q\bm{a}_{1}-\frac{q-1}{2}\bm{a}_{2}\right)\cdot\bm{b}_{1}&q\text{ odd}\end{cases}\right)=2\pi p\delta_{i,2},$ (22)
so that, indeed, in either case:
$\bm{a}_{i}\cdot\bm{A}(\bm{R}_{j})\in 2\pi\mathbb{Z}.$ (23)
Appendix B Threading flux
SI Affine-linear vector potential implementation
When dimensionally-reducing the lattice system to the cylinder, the charge and spin flux through the cross sections of the cylinder distinguish gauge-inequivalent systems. This additional flux can be implemented by adding a constant term to the vector potential:
$\bm{A}(\bm{r})\longrightarrow\bm{A}_{\phi}(\bm{r})=\bm{A}(\bm{r})+\phi\frac{\bm{b}_{2}}{2\pi L_{y}},$ (24)
where
$\phi=\phi_{q}+\frac{\sigma_{z}}{2}\phi_{s}.$ (25)
This modifies the hopping elements by
$t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}\longrightarrow t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}^{\phi}=t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}\exp\left(i\frac{\phi}{2\pi}\frac{\bm{b}_{2}\cdot\bm{\delta}}{L_{y}}\right).$ (26)
Since we made the choice $\bm{A}\propto\bm{b}_{2}$ for the original vector potential (in the absence of flux threading), then this is equivalent to a shift in the origin:
$\bm{A}_{\phi}(\bm{r})=\bm{A}(\bm{r}+\bm{X}_{\phi})=\bm{A}(\bm{r})+\frac{\Phi}{(2\pi)^{2}}(\bm{X}_{\phi}\cdot\bm{b}_{1})\bm{b}_{2},$ (27)
where
$\bm{X}_{\phi}=\frac{\phi\bm{a}_{1}}{\Phi L_{y}}.$ (28)
SII Twisted boundary condition implementation
For the purposes of numerical implementation, particularly in hybrid space where we exchange $y$ for $k_{y}$, it is convenient to instead implement flux threading via a twist in the boundary conditions. Starting with periodic boundary conditions $c^{\dagger}(\bm{r}+L_{y}\bm{a}_{2})=c^{\dagger}(\bm{r})$ for $\phi=0$, define a new set of operators
$c^{\dagger}_{\phi}(\bm{r})=c^{\dagger}(\bm{r})\exp(iy\phi/L_{y})=c^{\dagger}(\bm{r})\exp\left(i\phi\frac{\bm{b}_{2}\cdot\bm{r}}{2\pi L_{y}}\right).$ (29)
Then instead of implementing Eq. (24), we can instead replace every instance of $c$ with $c_{\phi}$. This does not affect terms that depend solely on charge and spin densities, e.g., the Hubbard interaction. However, it will modify the hoppings in a way that is equivalent to Eq. (24). Under the operator replacement, the hoppings are replaced by
$c^{\dagger}(\bm{r}+\bm{\delta})t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c(\bm{r})$ $\longrightarrow c_{\phi}^{\dagger}(\bm{r}+\bm{\delta})t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\phi}(\bm{r})$ (30)
$=c^{\dagger}(\bm{r}+\bm{\delta})\exp\left(i\phi\frac{\bm{b}_{2}\cdot(\bm{r}+\bm{\delta})}{2\pi L_{y}}\right)t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}\exp\left(-i\phi\frac{\bm{b}_{2}\cdot\bm{r}}{2\pi L_{y}}\right)c(\bm{r})$ (31)
$=c^{\dagger}(\bm{r}+\bm{\delta})\exp\left(i\frac{\phi}{2\pi}\frac{\bm{b}_{2}\cdot\bm{\delta}}{L_{y}}\right)t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c(\bm{r}),$ (32)
which is exactly equivalent to Eq. (26). To see that Eq. (29) corresponds to a twisting of the boundary condition on the fermion operators, we note that
$c^{\dagger}_{\phi}(\bm{r}+L_{y}\bm{a}_{2})=c^{\dagger}(\bm{r}+L_{y}\bm{a}_{2})\exp\left(i\phi\frac{\bm{b}_{2}\cdot(\bm{r}+L_{y}\bm{a}_{2})}{2\pi L_{y}}\right)=c_{\phi}^{\dagger}(\bm{r})\exp(i\phi).$ (33)
The utility of this alternative formulation is that it leads to a clean numerical implementation upon fourier transforming $y\to k_{y}$. In particular, recall that $(x,y)=(\bm{r}\cdot\bm{b}_{1}/2\pi,\bm{r}\cdot\bm{b}_{2}/2\pi)$ and define
$c_{\phi}^{\dagger}(x,k_{y})=\frac{1}{\sqrt{L_{y}}}\sum_{y}e^{-ik_{y}y}c_{\phi}^{\dagger}(x,y)\iff c_{\phi}^{\dagger}(x,y)=\frac{1}{\sqrt{L_{y}}}\sum_{k_{y}}e^{ik_{y}y}c_{\phi}^{\dagger}(x,k_{y}).$ (34)
Note that this is the Fourier convention of TeNPy. When we express the Hamiltonian in the hybrid $xk_{y}$ representation, we find that all that’s required is to shift the allowed values of $k_{y}.$ To see this, we impose the twisted boundary condition Eq. (33) on the Fourier-expanded operators:
$\frac{1}{\sqrt{L_{y}}}\sum_{k_{y}}e^{ik_{y}(y+L_{y})}c_{\theta}^{\dagger}(x,k_{y})=\frac{1}{\sqrt{L_{y}}}\sum_{k_{y}}e^{ik_{y}y}c_{\theta}^{\dagger}(x,k_{y})e^{i\phi},$ (35)
so that
$e^{ik_{y}L_{y}}=e^{i\phi}.$ (36)
The allowed values of momenta are therefore quantized to:
$k_{y}=\frac{2\pi}{L_{y}}\left(n+\frac{\phi_{q}}{2\pi}+\frac{\sigma_{z}\phi_{s}}{4\pi}\right),\qquad n\in\mathbb{Z},$ (37)
which are shifted relative to the $\phi=0$ case. When the original Hamiltonian $\hat{H}_{\phi=0}$ is written in hybrid space, some coefficients will depend on $k.$ The only affect of the replacement Eq. (29) is then for those coefficients to instead be evaluated on these shifted set of allowed values.
Appendix C Two-band hopping Hamiltonian
SI Hamiltonian at $(p,q)=(1,2)$
Now let us specialize to the case $(p,q)=(1,2)$, which corresponds to $\Phi_{\triangle}=\pi/2$. To specify the unit cells (which each contain two sites), we often write $x=\tau+2u$, where $u\in\mathbb{Z}$, while $\tau\in\{0,1\}$ specifies the even/odd rings. With periodic boundary conditions and before flux insertion, the hoppings Eq. (15) take the particular form
$t_{(x,y+1)\leftarrow(x,y)}$ $=t\exp\left(i\bm{A}\left(x\bm{a}_{1}+y\bm{a}_{2}+\frac{\bm{a}_{2}}{2}\right)\cdot\bm{a}_{2}\right)$
$=t\exp\left(\frac{i}{4\pi}\left(x\bm{a}_{1}+y\bm{a}_{2}+\frac{\bm{a}_{2}}{2}\right)\cdot\bm{b}_{1}\bm{b}_{2}\cdot\bm{a}_{2}\right)$
$=t\exp\left(i\pi x\right),$ (38)
$t_{(x+1,y)\leftarrow(x,y)}$ $=t\exp\left(i\bm{A}\left(\frac{\bm{a}_{1}}{2}+x\bm{a}_{1}+y\bm{a}_{2}\right)\cdot\bm{a}_{1}\right)$
$=t,$ (39)
$t_{(x+1,y-1)\leftarrow(x,y)}$ $=t\exp\left(i\bm{A}\left(\frac{\bm{a}_{1}-\bm{a}_{2}}{2}+x\bm{a}_{1}+y\bm{a}_{2}\right)\cdot(\bm{a}_{1}-\bm{a}_{2})\right)$
$=t\exp\left(-i\bm{A}\left(\frac{\bm{a}_{1}-\bm{a}_{2}}{2}+x\bm{a}_{1}+y\bm{a}_{2}\right)\cdot\bm{a}_{2}\right)$
$=t\exp\left(-i\pi(x+1/2)\right).$ (40)
We depict the resulting hopping phases in Fig. S1(a) for the lattice orientation where $\bm{a}_{2}=a\hat{y}$ is a nearest-neighbour bond; this becomes becomes the YC geometry when the system is placed on a cylinder under the identification $\bm{r}\sim\bm{r}+L_{y}\bm{a}_{2}$.
Figure S1: (a) Hopping phases for the $(p,q)=(1,2)$ two-band Hamiltonian corresponding to $\pi/2$ flux per triangle. The two black circles indicate the two sites in the magnetic unit cell. The phases indicated by the color bar are accrued when hopping upward, right-upward, and right-downward; the complex-conjugate phases are accrued when hopping against these directions. (b) Graphical depiction of the fluxes through various plaquettes upon placing the lattice on a cylinder. Every triangle on the surface of the infinite cylinder is penetrated by $\Phi_{\triangle}=\pi/2$ external flux. As a consequence of Gauss’s law, the external flux through adjacent rings—each of which is a polygonal surface bounded by a ring of length $L_{y}$—must alternate by $\pi$ when $L_{y}$ is odd, and must be uniform when $L_{y}$ is even.
SII Symmetries in the 2D plane
SII.1 Magnetic translations
Recall that we designed the vector potential so that the Hamiltonian would be invariant under $T_{y}$ translations, i.e., in the circumferential direction $\bm{a}_{2},$ defined as
$T_{y}c_{\bm{r}}^{\dagger}T_{y}^{-1}=c_{\bm{r}+\bm{a}_{2}}^{\dagger}.$ (41)
To see this, note that
$t_{\bm{r}+\bm{a}_{2}+\bm{\delta}\leftarrow\bm{r}+\bm{a}_{2}}$ $=t\exp\left(i\bm{A}\left(\bm{r}+\bm{a}_{2}+\bm{\delta}/2\right)\cdot\bm{\delta}\right)$ (42)
$=t\exp\left(i\bm{A}\left(\bm{r}+\bm{\delta}/2\right)\cdot\bm{\delta}\right)$ (43)
$=t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}},$ (44)
since $\bm{A}(\bm{a}_{2})=0$ by design. On the other hand, translation along $\bm{a}_{1}$ introduces an additional term:
$t_{\bm{r}+\bm{a}_{1}+\bm{\delta}\leftarrow\bm{r}+\bm{a}_{1}}$ $=t\exp\left(i\bm{A}\left(\bm{a}_{1}+\bm{r}+\bm{\delta}/2\right)\cdot\bm{\delta}\right)$ (45)
$=\exp\left(i\bm{A}\left(\bm{a}_{1}\right)\cdot\bm{\delta}\right)t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}$ (46)
$=\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{\delta}\right)t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}.$ (47)
This affects the hopping directions with some weight along $\bm{\delta}=\bm{a}_{2}$ and, by comparison to Eq. (28) above, corresponds to threading $\phi=\pi L_{y}$ through the cylinder. In the 2D plane geometry — and later on some even-$L_{y}$ geometries as we’ll show — the Hamiltonian can be restored by composing the bare $T_{x}$ with a large gauge transformation. In particular, define the magnetic translation operation by
$T_{x}^{\text{mag}}c_{\bm{r}}^{\dagger}(T_{x}^{\text{mag}})^{-1}=c_{\bm{r}+\bm{a}_{1}}^{\dagger}\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right)=c_{\bm{r}+\bm{a}_{1}}^{\dagger}(-1)^{y},$ (48)
so that
$T_{x}^{\text{mag}}\hat{h}(T_{x}^{\text{mag}})^{-1}$ $=T_{x}^{\text{mag}}\left(\sum c_{\bm{r}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\bm{r}}\right)(T_{x}^{\text{mag}})^{-1}$ (49)
$=\sum\exp\left(\frac{i}{2}\bm{b}_{2}\cdot(\bm{r}+\bm{\delta})\right)c_{\bm{r}+\bm{a}_{1}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\bm{r}+\bm{a}_{1}}\exp\left(-\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right)$ (50)
$=\sum c_{\bm{r}+\bm{a}_{1}+\bm{\delta}}^{\dagger}\left(\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{\delta}\right)t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}\right)c_{\bm{r}+\bm{a}_{1}}$ (51)
$=\sum c_{\bm{r}+\bm{a}_{1}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{a}_{1}+\bm{\delta}\leftarrow\bm{r}+\bm{a}_{1}}c_{\bm{r}+\bm{a}_{1}}$ (52)
$=\hat{h},$ (53)
where the second-last equality follows from Eq. (47) above.
SII.2 Particle-hole
Consider the naive definition of the particle-hole operation:
$\tilde{\mathcal{P}}c^{\dagger}(\bm{r})\tilde{\mathcal{P}}^{-1}=c(\bm{r}),\qquad\tilde{\mathcal{P}}i\tilde{\mathcal{P}}^{-1}=+i,$ (54)
where the condition specifies it is a linear (as opposed to anti-linear) unitary many-body operator. If 1 and 2 are an adjacent pair of sites, then particle-hole acts as
$\tilde{\mathcal{\mathcal{P}}}(t_{12}c_{1}^{\dagger}c_{2}+t_{12}^{*}c_{2}^{\dagger}c_{1})\tilde{\mathcal{\mathcal{P}}}^{-1}$ $=t_{12}c_{1}c_{2}^{\dagger}+t_{12}^{*}c_{2}c_{1}^{\dagger}$ (55)
$=-t_{12}c_{2}^{\dagger}c_{1}-t_{12}^{*}c_{1}^{\dagger}c_{2}$ (56)
$=-t_{12}^{*}c_{1}^{\dagger}c_{2}-t_{12}c_{2}^{\dagger}c_{1},$ (57)
so that it flips the Peierl’s phase and increases it by $\pi.$ On the triangular lattice, this modifies the flux through each triangle as follows:
$\Phi_{\triangle}\to-\Phi_{\triangle}+\pi.$ (58)
Requiring this operation to be invariant mod $2\pi,$ we find that we need
$\Phi_{\text{par}}=2\Phi_{\triangle}=\pi\left(2n+1\right),\qquad n\in\mathbb{Z}.$ (59)
This is consistent with the observation that the Hofstadter spectrum on the triangular lattice only has bands with opposite energy for these values of $\Phi_{\triangle}$. In particular, in this work we consider $\Phi_{\triangle}=\pi/2$. We find at each momentum $\bm{k}$ that both the energy and Berry curvature are exactly opposite, with the bands having Chern number ±$\pm 1.$
In general, the naive particle-hole must be composed with a gauge transformation in order to restore all the hopping elements. Given the particular form of the hoppings Eq. (SI) to Eq. (SI) for the two-band model, we obtain the following form of the particle-hole symmetry:
$\mathcal{P}c^{\dagger}(\bm{r})\mathcal{P}^{-1}=(-1)^{x+y}c^{\mathsf{T}}(\bm{r}),\qquad\mathcal{P}i\mathcal{P}^{-1}=+i,$ (60)
where $\mathsf{T}$ is the transposition operation on the row vector of electronic annihilation operators.
SII.3 Mirror composed with time-reversal
The orbital magnetic field breaks the time-reversal and mirror symmetries enjoyed by the $\Phi_{\triangle}\equiv 0$ (mod $\pi$) Hubbard models on the triangular lattice. However, the combination of mirror with time-reversal remains a symmetry and should not be spontaneously broken by the CSL or IQH phases. In particular, we focus on $M_{y}\mathcal{T}$ since this symmetry acts “on-site” (i.e., within each ring) when the system is placed on the cylinder. Recall that $M_{y}$ should send $\bm{r}\to M_{y}\bm{r}$ (flipping the Cartesian $y$-coordinate) while also mapping $\bm{S}\to(-S_{x},S_{y},-S_{z}).$ On the other hand, time-reversal should flip all components of spin. Their combination should therefore map $\bm{S}\to(S_{x},-S_{y},S_{z})$, which is easy to implement due to $\mathcal{T}$ being anti-unitary. We claim that the desired symmetry is
$(M_{y}\mathcal{T})c^{\dagger}(\bm{r})(M_{y}\mathcal{T})^{-1}=c^{\dagger}(M_{y}\bm{r})e^{i(-1)^{x}\pi/4},$ (61)
where the exponential factor is chosen to restore the phases of the hoppings. We confirm this by explicit calculation:
$(M_{y}\mathcal{T})\hat{h}(M_{y}\mathcal{T})^{-1}$ $=(M_{y}\mathcal{T})\left(\sum c_{\bm{r}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\bm{r}}\right)(M_{y}\mathcal{T})$ (62)
$=\sum c_{M_{y}(\bm{r}+\bm{\delta})}^{\dagger}e^{i(-1)^{\delta_{x}+x}\pi/4}t^{*}_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}e^{-i(-1)^{x}\pi/4}c_{M_{y}\bm{r}}$ (63)
$=\sum c_{\bm{r}+\bm{\delta}}^{\dagger}e^{i(-1)^{x}\left((-1)^{\delta_{x}}-1\right)\pi/4}t^{*}_{M_{y}(\bm{r}+\bm{\delta})\leftarrow M_{y}\bm{r}}c_{\bm{r}},$ (64)
which can be shown to be respected for Eqs. (SI) to (SI) above.
It is easy to verify that $M_{y}\mathcal{T}$ is not Kramers. However, it can be convenient to define a Kramers symmetry by performing an additional $\pi$ spin rotation of about the $y$ axis:
$(M_{y}\mathcal{T}^{\prime})c^{\dagger}(\bm{r})(M_{y}\mathcal{T}^{\prime})^{-1}=c^{\dagger}(M_{y}\bm{r})(i\sigma_{y})e^{i(-1)^{x}\pi/4}.$ (65)
One then has that
$(M_{y}\mathcal{T})^{2}c^{\dagger}(\bm{r})(M_{y}\mathcal{T})^{-2}$ $=+c^{\dagger}(\bm{r}),$ (66)
$(M_{y}\mathcal{T}^{\prime})^{2}c^{\dagger}(\bm{r})(M_{y}\mathcal{T}^{\prime})^{-2}$ $=-c^{\dagger}(\bm{r}).$ (67)
Furthermore, we note that
$(M_{y}\mathcal{T})\bm{S}(\bm{r})(M_{y}\mathcal{T})^{-1}$ $=\left(S_{x},-S_{y},S_{z}\right)_{M_{y}\bm{r}},$ (68)
$(M_{y}\mathcal{T}^{\prime})\bm{S}(\bm{r})(M_{y}\mathcal{T}^{\prime})^{-1}$ $=-\bm{S}(M_{y}\bm{r}).$ (69)
SIII Symmetries on the cylinder
To implement charge-flux threading, we work in the real-space formalism with periodic boundary conditions and vector potential modified according to Eq. (24). We then get
$t_{(x,y+1)\leftarrow(x,y)}^{\phi}$ $=t\exp\left(i\pi x\right)\exp\left(i\phi/L_{y}\right)$ (70)
$t_{(x+1,y)\leftarrow(x,y)}^{\phi}$ $=t$ (71)
$t_{(x+1,y-1)\leftarrow(x,y)}^{\phi}$ $=t\exp\left(-i\pi(x+1/2)\right)\exp\left(-i\phi/L_{y}\right),$ (72)
so that we pick up an additional phase of $\phi/L_{y}$ when $y$ increments by $1$ One finds that the symmetries of the system are sensitive not only to the parity of $L_{y}$, but also to the threaded charge and spin flux.
SIII.1 Magnetic translations
Figure S2: Density-density connected correlations for each of the four studied circumferences $L_{y}=3$ to $6$, with the Hubbard interaction tuned to the values at which the correlation length is maximal at the largest available bond dimension, namely $U(L_{y}=3)=11.125$, $U(L_{y}=4)=11.0$, $U(L_{y}=5)=11.6$, and $U(L_{y}=6)=11.4$, respectively. The even-circumference cylinders are threaded with $\Phi_{\text{ext}}=\pi/2$, as in the main text.
When $L_{y}$ is even, the Hamiltonian can be restored by composing the bare $T_{x}$ with a large gauge transformation. In contrast, by Fig. S1(b), we know that the effect of translations on the fluxes through loops cannot be gauged away when $L_{y}$ is odd. This manifests in the fact that $T_{x}^{\text{mag}}$ [see Eq. (48)] is ill-defined, which we see by imposing periodic boundary conditions and comparing $\bm{r}+L_{y}\bm{a}_{2}$ with $\bm{r}$; note that
$\exp\left(\frac{i}{2}\bm{b}_{2}\cdot(\bm{r}+L_{y}\bm{a}_{2})\right)=\exp\left(i\pi L_{y}\right)\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right)\neq\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right).$ (73)
SIII.2 Particle-hole
Now, let’s fix $\Phi_{\text{par}}=2\Phi_{\triangle}=+\pi$ and interrogate whether placing the system on a YC cylinder changes this conclusion. Focusing on a particular ring $r$ around the cylinder with flux $\Phi_{r},$ we note that
$\tilde{\mathcal{\mathcal{P}}}:\Phi_{r}\mapsto\begin{cases}-\Phi_{r}&L_{y}\text{ even},\\ -\Phi_{r}+\pi&L_{y}\text{ odd}.\end{cases}$ (74)
Therefore, $\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to a gauge transformation) when
$\Phi_{r}\in\begin{cases}\pi\mathbb{Z}&L_{y}\text{ even},\\ \frac{\pi}{2}+\pi\mathbb{Z}&L_{y}\text{ odd},\end{cases}$ (75)
for all rings $r.$
When $L_{y}$ is even, Gauss’s law dictates that the flux $\Phi_{r_{1,2}}$ through two adjancent rings $r_{1,2}$ must be equal — note that this allows $T_{x}$ to be a symmetry (up to a gauge transformation). Therefore, possibly after threading $\phi$ external flux through each ring of the cylinder, it is possible to tune all the $\Phi_{r}$ simultaneously (to 00 or $\pi$) so that $\tilde{\mathcal{\mathcal{P}}}$ is respected. Only in those cases is $T_{x}\tilde{\mathcal{\mathcal{P}}}$ also a symmetry (up to gauge).
On the other hand, when $L_{y}$ is odd, Gauss’s law instead dictates that
$\Phi_{r_{2}}-\Phi_{r_{1}}=\pi\qquad\text{mod }2\pi.$ (76)
This always rules out $T_{x}$ as a symmetry. However, so long as $\Phi_{r_{1}}$ is $\pi/2$ or $-\pi/2,$ both Eqs. (75) and (76) can be respected, so that $\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to gauge). However, odd $L_{y}$ is special in that if we tune $\Phi_{r_{1}}$ to 0 or $\pi,$ then $T_{x}\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to gauge) even though neither $T_{x}$ nor $\tilde{\mathcal{\mathcal{P}}}$ are themselves symmetries.
Appendix D Observable and Correlation Functions
SI Heisenberg Operator in terms of Fermions
If the sites are spinless FermionSites and up and down spins are encoded in adjacent FermionSites, we can compute spin correlations indirectly using the fermion creation and annihilation operators. Since the spin correlation operator $\vec{S}_{i}\cdot\vec{S}_{j}$ involves the product of two spin operators, we need to express it in terms of the fermion operators. In a system with spinless FermionSites, we can define the spin operators as:
$S_{i}^{+}=c_{i\uparrow}^{\dagger}c_{i\downarrow},\quad S_{i}^{-}=c_{i\downarrow}^{\dagger}c_{i\uparrow},\quad S_{i}^{z}=\frac{1}{2}(c_{i\uparrow}^{\dagger}c_{i\uparrow}-c_{i\downarrow}^{\dagger}c_{i\downarrow}),$ (77)
where $c_{i\uparrow}^{\dagger}$ creates an up-spin fermion on site $i$ and $c_{i\downarrow}^{\dagger}$ creates a down-spin fermion on site $i$. The spin correlation operator $\vec{S}_{i}\cdot\vec{S}_{j}$ can be written in terms of these fermion operators as:
$\vec{S}_{i}\cdot\vec{S}_{j}=\frac{1}{2}\left(c_{i\uparrow}^{\dagger}c_{i\downarrow}c_{j\downarrow}^{\dagger}c_{j\uparrow}+c_{i\downarrow}^{\dagger}c_{i\uparrow}c_{j\uparrow}^{\dagger}c_{j\downarrow}\right)+S_{i}^{z}S_{j}^{z}.$ (78)
SII Staggered Heisenberg Order Parameter
Here we express the staggered-Heisenberg order parameter in the hybrid $(x,k_{y})$ operator representation, where $x$ indexes cylinder rings. Because of translation invariance around the cylinder, it is useful to define the $y$-averaged inter-ring NN Heisenberg operator
$\mathcal{O}_{\text{H}}(x)=\frac{1}{L_{y}}\sum_{y}\sum_{a=x,y,z}S^{a}(x,y)S^{a}(x+1,y),$ (79)
where
$S^{a}(x,y)=c^{\dagger}(x,y)\frac{\sigma^{a}}{2}c(x,y)=\frac{1}{L_{y}}\sum_{k,k^{\prime}}e^{i(k-k^{\prime})y}c^{\dagger}(x,k)\frac{\sigma^{a}}{2}c(x,k^{\prime}).$ (80)
Since both the IQH and CSL ground states exhibit a two-ring translation invariance (i.e., are invariant under $(T_{x})^{2}$ in addition to $T_{y}$) then we can measure the local order parameter
$\mathcal{O}^{\delta\text{Heis}}_{0}=\mathcal{O}_{\text{H}}(1)-\mathcal{O}_{\text{H}}(0).$ (81)
In general, the hybrid representation reads
$\mathcal{O}_{\text{H}}(x)$ $=\frac{1}{L_{y}}\sum_{y}\sum_{a=x,y,z}S^{a}(x,y)S^{a}(x+1,y)$
$=\frac{1}{L_{y}^{2}}\sum_{a=x,y,z}\sum_{k,k^{\prime}}\sum_{p,p^{\prime}}\left(\frac{1}{L_{y}}\sum_{y}e^{i(k-k^{\prime}+p-p^{\prime})y}\right)c^{\dagger}(x,k)\frac{\sigma^{a}}{2}c(x,k^{\prime})\cdot c^{\dagger}(x+1,p)\frac{\sigma^{a}}{2}c(x+1,p^{\prime})$
$=\frac{1}{L_{y}^{2}}\sum_{a=x,y,z}\sum_{k,k^{\prime}}\sum_{p,p^{\prime}}c^{\dagger}(x,k)\frac{\sigma^{a}}{2}c(x,k^{\prime})\cdot c^{\dagger}(x+1,p)\frac{\sigma^{a}}{2}c(x+1,k-k^{\prime}+p).$
Appendix E Correlation Lengths
Figure S3: Correlation length data for even-circumference geometries. Left panel: Illustration of the external magnetic flux penetrating each loop of the infinite cylinder, namely no flux, $\Phi=0$. Center panels: Maximum correlation length at $L_{y}=4$ in the $Q=0e$ charge sector [blue] for a range of bond dimensions increasing with increasing darkness; and in the $Q=1e$ charge sector [green]. Right panel: The same for $L_{y}=6$.
Figure S4: Correlation length data for even-circumference geometries. Left panel: Illustration of the external magnetic flux penetrating each loop of the infinite cylinder, which alternates between ±$\pm\pi/2$, corresponding to threaded flux $\Phi_{\text{ext}}=\pi/2$ in our conventions. Center panels: Maximum correlation length at $L_{y}=3$ in the $Q=0e$ charge sector [blue] for a range of bond dimensions increasing with increasing darkness; and in the $Q=1e$ charge sector [green]. Right panel: The same for $L_{y}=5$.
Here we briefly comment on additional data concerning the correlation lengths for geometries not displayed in the text. In particular, Fig. S3 shows correlation length data for the $L_{y}=4$ and $6$ cylinders which are threaded by no external flux. In particular, $L_{y}=6$ reproduces the behavior presented in Kuhlenkamp et al., [11]. We remark that the correlation lengths are quantitatively smaller than those at $\Phi_{\text{ext}}$ external flux, which are shown in the main text. We similarly show data for the $L_{y}=3$ and $5$ cylinders threaded by $\Phi_{\text{ext}}=\pi/2$ external flux. Unlike at $\Phi_{\text{ext}}=0$ in the main text, the flux breaks the glide-PH symmetry, and thus we see no diverging or quantitatively large correlation length (as is consistent with no spontaneous symmetry breaking).
Appendix F Correlation Length Extrapolation
Figure S5: (a) Extrapolated inverse correlation length in the charge-neutral sector at $L_{y}=4$ and $\Phi_{\text{ext}}=\pi/2$ obtained following the method in Ref. [52], as detailed in Sec. F, for a range of Hubbard strengths [different colors] and for a sequence of increasing bond dimensions [same-colored points moving sequentially to the left]. The extrapolated inverse correlation length saturates at a minimum value exceeding $0.15$; the correlation length is equivalently bounded above by $6.6$ ring spacings. (b) The same for $L_{y}=5$ and $\Phi_{\text{ext}}=0$. In this case, the extrapolated inverse correlation length reaches and descends beyond 0, signaling the need for larger bond dimensions to accurately estimate large correlation lengths.
To obtain the extrapolated correlation lengths shown in the main text, we employ a recently-developed method for extrapolating correlation lengths in DMRG [52, 59]. Namely, we compute several of the largest eigenvalues of the transfer matrix $TX_{i}=\lambda_{i}X_{i}$ in the electric charge $Q_{E}=1$ sector for each $k_{y}$, and select those from the same “excitation branch” [52], namely carrying a consistent complex phase. These are related to the correlation lengths as
$\epsilon_{i}=1/\xi_{i}=-\log\left|\lambda_{i}\right|$ (82)
with $\epsilon_{i}\leq\epsilon_{i+1}$. We report $\xi$ in units of cylinder rings, or equivalently $u=a\sqrt{3}/2$ where $a$ is the lattice spacing of the underlying triangular lattice. Correspondingly, $\epsilon$ is reported in units of $u^{-1}$. For a system that is gapped in a charge sector $q=(Q,S_{z},k_{y})$, both $\epsilon_{1}^{q},\epsilon_{2}^{q}$ converge to the same infinite value as $\chi\to\infty$, namely the finite inverse correlation length. As a result, the quantity
$\delta^{q}=\epsilon_{2}^{q}-\epsilon_{1}^{q}$ (83)
vanishes at $\chi\to\infty$, serving as a good “scaling variable” that measures the deviation from convergence. To estimate $\xi(\chi\to\infty)$, we therefore compute $\epsilon(\delta)$ at each available $\epsilon(\chi),\delta(\chi)$ and extrapolate to $\delta=0$ with a polynomial fit, namely
$\epsilon(\delta)=a+b\delta^{c}.$ (84)
This extrapolation is shown in Fig. S5 for both $L_{y}=4$ at $\Phi_{\mathrm{ext}}=\pi/2$ and $L_{y}=5$ at $\Phi_{\mathrm{ext}}=0$, yielding the extrapolated values shown in the main text. In both cases, the correlation length is computed in the $(Q,S_{z},k_{y})=\bm{0}$, which hosts largest correlation length.
Appendix G Density-density correlations at all circumferences
Figure S6: Density-density connected correlations for each of the four studied circumferences $L_{y}=3$ to $6$, with the Hubbard interaction tuned to the values at which the correlation length is maximal at the largest available bond dimension, namely $U(L_{y}=3)=11.125$, $U(L_{y}=4)=11.0$, $U(L_{y}=5)=11.6$, and $U(L_{y}=6)=11.4$, respectively. The even-circumference cylinders are threaded with $\Phi_{\text{ext}}=\pi/2$, as in the main text.
In Fig. S6 we show the ring-averaged density-density connected correlations, defined as $\langle n_{\text{av}}(x)n_{\text{av}}(0)\rangle,$ where
$n_{\text{av}}(x)$ $=\frac{1}{L_{y}}\sum_{y}c^{\dagger}(x,y)c(x,y)=\frac{1}{L_{y}}\sum_{k}c^{\dagger}(x,k)c(x,k).$ (85)
Panels (a,c), which correspond to the odd circumferences $L_{y}=3,5$ respectively, exhibit apparently algebraic correlations at long distances. On the other hand, panels (b,d) exhibit exponentially-decaying correlations that nonetheless feature a near-algebraic plateau, over a range of $x$ less than their respective correlation lengths.

### v2

@@
 Abstract
-Recent advancements in moiré engineering motivate study of the behavior of strongly-correlated electrons subject to substantial orbital magnetic flux.
-We investigate the triangular lattice Hofstadter-Hubbard model at one-quarter flux quantum per plaquette and a density of one electron per site, where geometric frustration has been argued to stabilize a chiral spin liquid phase intermediate between the weak-coupling integer quantum Hall and strong-coupling 120 antiferromagnetic phases.
-In this work, we use Density Matrix Renormalization Group methods and analytical arguments to analyze the compactification of the Hofstadter-Hubbard model to cylinders of finite radius.
-We introduce a glide particle-hole symmetry operation which for odd-circumference cylinders, we show, is spontaneously broken at the quantum Hall to spin liquid transition.
-We further demonstrate that the transition is associated with a diverging correlation length of a charge-neutral operator.
-For even-circumference cylinders the transition is associated with a dramatic quantitative enhancement in the correlation length upon threading external magnetic flux.
-Altogether, we argue that the 2+1D CSL-IQH transition is in fact continuous and features critical correlations of the charge density and other spin rotationally-invariant observables.
-Introduction—The Kalmeyer-Laughlin chiral spin liquid (CSL) is one of the earliest and most important examples of topological order [1, 2, 3] and has been extensively studied in spin systems, both numerically [4] and through partonic mean-field approaches [5, 6, 7].
-Recently, considerable efforts have been devoted to identifying electronic settings in which the CSL may emerge to connect more directly to solid state experiments, in which electrons—rather than spins—are the fundamental degrees of freedom and may be far from their Mott limit.
-These systems enable the exploration of the Mott transition of strongly-interacting electrons in strong magnetic fields, a regime that is becoming experimentally accessible both at non-zero [8, 9] and zero Zeeman coupling [10, 11].
-They also set the stage for the study of charge fluctuation-driven transitions out of the CSL phase into other phases such as superconductors [12], exotic charge-density waves [13], or quantum Hall states [14, 11].
-Figure 1: (a) Upper panel: The Hofstadter-Hubbard model on the 2D triangular lattice, with magnetic translations $T_{x,y}$, on-site Hubbard $U$ interaction, and flux per triangular plaquette $\Phi_{\triangle}$ indicated.
-Lower left panel: “Hofstadter butterfly” representation of spectrum of non-interacting model in the plane of single-particle energy $E$ and flux per plaquette $\Phi_{\triangle}$.
-$\Phi_{\triangle}=\pi/2$ is the unique flux for which the Hamiltonian enjoys a doubled “magnetic” unit cell and a particle-hole symmetry $\mathcal{P}$ relating opposite $C=\pm 1$ bands [lower right panel].
-(b) A finite segment of the infinite-length, circumference-$L_{y}$ cylinder geometry, threaded by $\Phi_{\text{ext}}$ external magnetic flux.
-With the chosen “YC” cylindrical geometry, the translation $T_{y}$ is orthogonal to the cylinder axis and $x\in\mathbb{Z}$ indexes rings.
-(c) Expected 2+1D phase diagram at half-filling.
-As interactions increase, the spin-degenerate integer quantum Hall (IQH) ground state undergoes a quantum phase transition into a Kalmeyer-Laughlin chiral spin liquid (CSL).
-By studying both even and odd finite-circumference cylinders, we shed light on the existence and nature of this transition.
-A model electronic system thought to host both conventional and CSL electronic phases is the triangular Hofstadter-Hubbard model (see Fig.
-1a) in the regime of strong orbital magnetic flux, where the CSL may be favored energetically by both geometric frustration [15, 16] and the explicit breaking of time-reversal symmetry [17, 7].
-The Hofstadter-Hubbard model is parametrized by an on-site interaction $U$ and a nearest neighbor hopping with Peierls phases corresponding to a flux of $\Phi_{\triangle}$ per elementary plaquette.
-When $\Phi_{\triangle}=\pi/2$ per triangle, this system realizes a spin-unpolarized integer quantum Hall (IQH) insulator at small values of the Hubbard interaction, with $\sigma_{xy}=2e^{2}/h$, obtained by filling the lower topological sub-band in both spin sectors.
-This state possesses a gap to spin and charge excitations, is topological and has a spin-degenerate charged edge mode.
-At very large $U$ and density of one electron per site, the state is a Mott insulator: the leading term in the $t/U$ expansion around the infinite-$U$ Mott limit shows that the low-energy physics is governed by the conventional antiferromagnetic nearest-neighbor $S=1/2$ Heisenberg model with $J_{H}\sim t^{2}/U$, which has $120^{\circ}$ antiferromagnetic order at $T=0$, so that the ground state is non-topological, with fully-gapped charge excitations and soft (Goldstone) spin excitations.
-Carrying out the $t/U$ expansion to higher order suggests the existence of a CSL phase intermediate between the Mott/AF and IQH phases: the next correction is a third-order chiral term $S\cdot S\times S$ with coefficient $J_{\triangle}\sim t^{3}\sin\Phi_{\triangle}/U^{2}$.
-In the presence of the magnetic flux this term explicitly breaks both time-reversal and parity symmetry [17, 7].
-Numerical studies of the $J_{H}$-$J_{\triangle}$ spin Hamiltonian have indeed revealed a CSL ground state for large enough $J_{\triangle}/J_{H}$ [18, 19], although at these values the accuracy of truncating the $t/U$ expansion at third order may be questioned.
-Very recently Kuhlenkamp et al.
-presented cylinder iDMRG evidence of an intermediate CSL phase for $U\gtrsim 10t$ at fluxes $\Phi_{\triangle}=\pi/3,\pi/2$ [11].
-The CSL phase was identified by the absence of charge pumping as flux was threaded through the cylinder and the presence of appropriate degeneracies in the entanglement spectrum.
-Supposing one accepts the existence of an intermediate CSL phase, the question of the nature of the IQH-CSL transition arises.
-A parton theory in which one decomposes the electron into a bosonic chargon $b$ carrying the charge and a fermionic spinon $f$ carrying the spin index, $c_{\sigma}=bf_{\sigma}$ has been proposed [14, 12].
-This decomposition requires an emergent gauge field $a$, corresponding to the relative phase between the partons, under which they carry opposite unit $U(1)$ charge.
-In both the IQH and CSL phases the spinon occupies both $C=1$ magnetic sub-bands descending from the electronic bandstructure, while at the electronic IQH-CSL phase transition the chargon and the gauge field together undergo a superconductor-Mott phase transition [12].
-The dynamical gauge field $a$ is responsible not only for producing the correct topological response on both sides of the transition, but also for gapping out the charge sector in the IQH phase through the Anderson-Higgs mechanism [20, 21], and allowing for spinon deconfinement in the CSL [22, 23, 24].
-The parton representation and related slave-particle numerical studies on honeycomb and square lattices  [14, 25] have provided a physically intuitive, yet incomplete mean-field picture of the putative continuous transition.
-The iDMRG investigation of Ref.
-[11] revealed the IQH-CSL transition only as a crossover indicated by a weak maximum in a correlation length.
-Further investigation of the CSL phase and the IQH-CSL transition is therefore needed.
-In this work we use a compactification to an infinite-length, finite-radius cylinder [see Fig.
-1(b)] in combination with a symmetry analysis with large-scale iDMRG simulations [26, 27] on cylinders of circumference $L_{y}=3$ to $6$ to characterize the nature of the IQH and CSL states, as well as the transition between them.
-Our symmetry analysis is based on an operator that is a combination of particle-hole interchange and translation, which we show explicitly is spontaneously broken for interactions above a critical interaction strength for odd-circumference cylinders.
-We further show that the broken symmetry occurs together with the degeneracy pattern in the entanglement spectrum expected for a CSL phase.
-We present numerical evidence that this transition is in fact continuous and argue that in the limit of infinite cylinder radius, this transition evolves into a continuous 2+1D IQH-CSL quantum phase transition featuring critical fluctuations in the charge density and other spin rotationally-invariant observables.
-Figure 2: CSL symmetry-breaking at odd circumference.
-(a) Graphical illustration of the action of the particle-hole symmetry operator $\mathcal{P}$ which transforms the gauge flux through a ring of an odd-$L_{y}$ cylinder, from $\phi$ to $\pi-\phi$.
-(b) Upper panel: In an odd-$L_{y}$ CSL ground state, the alternating pattern $+\pi/2,-\pi/2,\dots$ of these gauge fluxes spontaneously breaks $\mathcal{P}T_{x}$ symmetry, whose action is equivalent to the threading of $\pi$ gauge flux.
-Lower panel: the low-lying entanglement eigenvalues at cuts between two pairs of rings, which not only exhibit the approximate degeneracy expected for the CSL at each $S_{z}$ [see circled numerals], but also transform as half-integer spin representations only at every other cut, signalling SSB of $\mathcal{P}T_{x}$.
-(c) In contrast, the internal gauge flux configuration [upper panel] and entanglement spectrum [lower panel] of the IQH state are consistent with it being a unique, symmetry-preserving ground state.
-Model, symmetry operator and cylinder compactification—We study the Hofstadter-Hubbard model on the triangular lattice, defined pictorially in Fig.
+Recent advances in moiré engineering motivate the study of lattice models of strongly-correlated electrons subjected to substantial orbital magnetic flux.
+We analyze the triangular lattice Hofstadter-Hubbard model at one-quarter flux quantum per plaquette and a density of one electron per site, where a chiral spin liquid phase may exist between weak-coupling integer quantum Hall and strong-coupling 120 antiferromagnetic phases.
+We use matrix product state methods and analytical arguments to investigate this model compactified to cylinders of finite circumference.
+We uncover a glide particle-hole symmetry operation which, we argue, is spontaneously broken at the quantum Hall to spin liquid transition on odd-circumference cylinders.
+We numerically verify the spontaneous symmetry breaking and further demonstrate that this transition is associated with algebraic long-range correlations of various spin-singlet, charge-neutral operators.
+For even-circumference cylinders, the transition becomes a crossover associated with a large correlation length that grows substantially with circumference.
+Our findings suggest that in the two-dimensional limit, the transition to a chiral spin liquid phase is continuous and features critical fluctuations of the current.
+Introduction—The Kalmeyer-Laughlin chiral spin liquid (CSL) is one of the earliest and most important examples of a topologically-ordered quantum state [1, 2, 3] and has been extensively studied in the context of spin models [4, 5, 6, 7].
+Recently, considerable effort has been devoted to identifying electronic settings in which the CSL may emerge [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
+These studies set the stage for the investigation of charge fluctuation-driven transitions out of the CSL phase into other phases such as metals [8], superconductors [20], exotic charge-density waves [21], and quantum Hall states [22, 19].
+Figure 1: (a) Upper panel: The triangular Hofstadter-Hubbard model, with magnetic translations $T_{x,y}$, flux per triangle $\Phi_{\triangle}$, and Hubbard interaction $U$ indicated.
+Lower left panel: “butterfly” representation of non-interacting spectrum in the plane of $\Phi_{\triangle}$ and energy $E$.
+The bands at $\Phi_{\triangle}=\pi/2$ enjoy a doubled “magnetic” unit cell and particle-hole symmetry $\mathcal{P}$ relating opposite $C=\pm 1$ bands [lower right panel].
+(b) Finite segment of the circumference-$L_{y}$ cylinder threaded by $\Phi_{\text{ext}}$ external flux, where $T_{y}$ translates around the cylinder and $x\in\mathbb{Z}$ indexes rings.
+(c) 2+1D phase diagram at a density of $n=1$ electron per site as a function of $U/t$ showing the integer quantum Hall (IQH) and chiral spin liquid (CSL) phases, with their transition marked by a star.
+The triangular Hofstadter-Hubbard model is thought to host both conventional and CSL electronic phases in the regime of strong orbital magnetic flux [23, 24, 6, 25, 26, 27, 28].
+It is parameterized by an on-site interaction $U$ and hoppings of magnitude $t$ and phases consistent with a magnetic flux of $\Phi_{\triangle}$ per triangular plaquette (see Fig.
+1a).
+When $\Phi_{\triangle}=\pi/2$, this system realizes a spin-singlet integer quantum Hall (IQH) insulator at small $U$ and density of one electron per site, with $\sigma_{xy}=2e^{2}/h$.
+At very large $U$, the state is a non-topological Mott insulator with a large charge gap, $120^{\circ}$ antiferromagnetic (AF) order, and gapless spin excitations described by a nearest-neighbor Heisenberg exchange of magnitude $\sim t^{2}/U$ [25, 26, 27, 28].
+A qualitative argument for the existence of an intermediate CSL phase is that the leading correction to the large-$U$ effective Heisenberg spin Hamiltonian is a chiral term $S\cdot S\times S$ with coefficient $J_{\triangle}\sim t^{3}\sin\Phi_{\triangle}/U^{2}$, which explicitly breaks both time-reversal and parity symmetry and favors a CSL state [24, 6, 29, 25, 26, 27, 28].
+Very recently, Kuhlenkamp et al.
+presented cylinder iDMRG evidence of this intermediate CSL phase starting from $U\gtrsim 8.5t$ at flux $\Phi_{\triangle}=\pi/3$ and $U\gtrsim 10.5t$ at $\Phi_{\triangle}=\pi/2$ [19].
+Figure 2: CSL symmetry breaking at odd circumference.
+(a) The particle-hole operation $\mathcal{P}$ transforms the internal gauge flux through an odd-$L_{y}$ cylinder ring from $\phi$ to $\pi-\phi$.
+(b) Upper panel: In an odd-$L_{y}$ CSL ground state, the alternating pattern $(\pi/2,-\pi/2,\dots)$ of these internal fluxes at strong coupling spontaneously breaks $\mathcal{P}T_{x}$ symmetry, whose action is then equivalent to threading $\pi$ internal flux.
+Lower panel: the low-lying entanglement eigenvalues on the $L_{y}=5$ cylinder at spatial cuts between adjacent pairs of rings, which not only exhibit the approximate degeneracy expected for the CSL at each fixed $S_{z}$ [circled numerals], but also signal spontaneous breaking of $\mathcal{P}T_{x}$.
+(c) The internal flux configuration [upper panel] and low-lying entanglement eigenvalues [lower panel] of the IQH at $L_{y}=5$.
+The entanglement spectra at the two spatial cuts are equivalent.
+The nature of the IQH-CSL transition, however, is not well understood.
+Parton mean field theories have been proposed for square lattice and honeycomb systems [22, 30].
+Beyond this, there has been very little direct investigation into the nature of the putative critical point from either theory or numerics.
+In particular, the iDMRG investigation of Ref.
+[19] unearthed the IQH-CSL transition only as a crossover indicated by a weak maximum in a correlation length.
+In this work, we use a compactification to infinite-length, finite-circumference cylinders (see Fig.
+1b) in combination with a symmetry analysis and large-scale iDMRG simulations [31, 32, 33] to characterize the IQH and CSL states, as well as the transition between them.
+We show that a combination of particle-hole interchange and translation is spontaneously broken on odd-circumference cylinders in the CSL phase and present evidence that the IQH-CSL transition is in fact continuous in these geometries, with no intervening phase.
+We argue that the critical fluctuations are spin-singlet, with both spin and electron correlation functions decaying exponentially with distance.
+Considering also even-circumference cylinders, we argue that in 2+1D the IQH and CSL phases are separated by a continuous quantum phase transition that includes critical current fluctuations.
+Figure 3: Evidence for the IQH-CSL continuous phase transition.
+(a) Upper panel: Sketch of two cylinder rings, illustrating the $\mathbb{Z}_{2}$ order parameter $\mathcal{O}_{0}$ defined as the total circumferential current operator [orange arrows].
+Lower panel: $\langle\mathcal{O}_{0}\rangle$ versus $U/t$ for the odd-circumference cylinders, with $\langle\mathcal{O}_{0}\rangle\neq 0$ implying $\mathcal{P}T_{x}$ is spontaneously broken.
+(b-d) MPS transfer matrix correlation length in the $(Q,S_{z})=0$ charge sector for various MPS bond dimensions [increasing light to dark] and extrapolated to infinite bond dimension for $L_{y}=4$ [gray points]; the insets specify the external magnetic flux through the cylinder rings.
+Model, particle-hole symmetry and compactification—We study the Hofstadter-Hubbard model on the triangular lattice, defined pictorially in Fig.
 1(a).
-In this model, electrons hop between nearest-neighboring sites and experience an Aharonov-Bohm phase of $\Phi_{\triangle}$ upon orbiting a triangular plaquette.
-Note that this model focuses on the consequences of the orbital motion, and sets the Zeeman coupling to zero.
-We fix $\Phi_{\triangle}=\pi/2$, which endows the Hamiltonian with a unitary particle-hole (PH) symmetry $\mathcal{P}$ acting as [28]
+In this model, electrons hop between nearest-neighboring sites and experience an Aharonov-Bohm phase of $\exp(i\Phi_{\triangle})$ upon orbiting a triangular plaquette (see App.
+A for details).
+We focus on the consequences of the orbital motion, setting the Zeeman coupling to zero [18, 19].
+We fix $\Phi_{\triangle}=\pi/2$, which endows the Hamiltonian with a unitary particle-hole (PH) symmetry $\mathcal{P}$ chosen to act as
 $\mathcal{P}c^{\dagger}_{x,y}\mathcal{P}^{-1}=(-1)^{x+y}(i\sigma_{y})c_{x,y},\quad\mathcal{P}i\mathcal{P}^{-1}=+i,$ (1)
-on fermionic operators, where we have attached a conventional spin rotation $\sigma_{y}$ to the naive definition of particle-hole transformation in order for $\mathcal{P}$ to leave the spin operator $\bm{S}=\frac{1}{2}c^{\dagger}\bm{\sigma}c$ invariant (see SM for details [29]).
-The flux induces a magnetic translation algebra $T_{x}T_{y}=-T_{y}T_{x}$ that we represent in a gauge where
-$T_{x}c^{\dagger}_{x,y}T_{x}^{-1}=(-1)^{y}c^{\dagger}_{x+1,y},\quad T_{y}c^{\dagger}_{x,y}T_{y}^{-1}=c^{\dagger}_{x,y+1}.$ (2)
-The commuting translations $(T_{x})^{2},T_{y}$ define a larger “magnetic” unit cell, whose resulting two sub-bands are related by $\mathcal{P}$ and carry opposite Chern number $C=\pm 1$.
-We fix the filling to $n=1$ electron per lattice site, which in the absence of interactions corresponds to full filling of the lowest spin-degenerate magnetic sub-band.
-We now present the dimensional reduction of the problem to a cylinder.
-The key results of our study rely on the different symmetry breaking patterns that occur for even and odd circumferences $L_{y}$, which stem from generic properties of magnetic fields (both external and internal) in the chosen geometry.
-To illustrate how this arises, let us first consider the closed “tin can”-shaped surface specified by a pair of adjacent rings at $x$ and $x+1$.
-This surface is the union of $2L_{y}$ triangles on the outside of the cylinder, as well as an $L_{y}$-sided polygonal surface at $x$ and another at $x+1$ whose edges are the rings.
-Since $\pi/2$ external flux penetrates each outer triangle, Gauss’s law requires that the external magnetic flux through the two adjacent rings differ by $\pi L_{y}$ [30, 31].
-These ring fluxes are therefore uniform when $L_{y}$ is even, but are staggered when $L_{y}$ is odd, which explicitly breaks $T_{x}$ symmetry.
-Let us momentarily focus on the odd case and fix the sequence of ring fluxes to alternate between 00 and $\pi$.
-In this case, the system nonetheless respects a combined translation-like symmetry, namely the product of $\mathcal{P}$ and $T_{x}$—which we refer to as PH-glide—in spite of each being individually broken by the compactification.
-The IQH state, being adiabatically connected to the (unique) ground state of the non-interacting Hamiltonian, is invariant under PH-glide symmetry.
-We argue, however, that the CSL phase must break $\mathcal{P}T_{x}$ spontaneously when $L_{y}$ is odd.
-To see this, note that the CSL is a Mott insulating phase.
-Thus, we may consider the limit where charge fluctuations are frozen out, e.g.
-by increasing $U\to\infty$ while holding the three-spin chiral term at a fixed favorable strength.
-Since the particle-hole operation is transparent to the spin operator, $\mathcal{P}\bm{S}\mathcal{P}^{-1}=\bm{S}$, this results in a spin model with a pure translation symmetry $t_{x}$.00footnotetext: We note a similar observation in a simpler 1+1D model [32].
-Since each ring contains an odd number of spin-$1/2$ moments, the Lieb-Schulz-Mattis theorem requires that any gapped ground state be at least two-fold degenerate [33], e.g., by spontaneously breaking $t_{x}$.
-One way to understand the translation symmetry breaking of the CSL, in particular, is by leveraging its anyonic content.
-Namely, the properties of the CSL under translations are captured by viewing it has hosting a spinon in each unit cell of the triangular lattice [34, 35, 36].
-Because these spinons have semionic statistics, the action of $t_{x}$ translation—which moves $L_{y}$ many spinons across any given cut—necessarily permutes the two minimally-entangled CSL ground states [37] when $L_{y}$ is odd [38, 39].
-In other words, the only $t_{x}$-symmetric ground states are long-range-entangled CSL “cat” states, implying that translation is spontaneously broken [40].
-Figure 3: Evidence for the IQH-CSL phase transition.
-(a) Upper panel: Sketch of three rings of cylinder at indicated $x$ values, with solid and dashed lines showing along which staggered-Heisenberg order parameter $\mathcal{O}_{\delta\mathrm{Heis}}$ [defined in Eq.
-(3)] changes under glide particle-hole symmetry $\mathcal{P}T_{x}$ on odd-circumference cylinders.
-Lower panel: Magnitude of the staggered Heisenberg order parameter plotted against $U/t$ for cylinder widths indicated and with $\mathcal{P}T_{x}$ preserving and broken phases indicated by the check and cross symbols.
-Consistent with the recovery of this symmetry at the putative 2+1D critical point, the magnitude of the order parameter [bottom left] decreases substantially from $L_{y}=3$ to $5$.
-(b-d) Correlation length in the $(Q,S_{z},k_{y})=\bm{0}$ charge sector, plotted against interaction strength, for a range of MPS bond dimensions [increasing light to dark] and extrapolated to infinite bond dimension [gray points]; the insets specify the external flux threading each of the cylinder rings.
-(b) At circumference $L_{y}=5$, the $\mathcal{P}T_{x}$ symmetry-breaking is associated with a diverging correlation length.
-(c) While both the IQH and CSL states are fully-symmetric on even-$L_{y}$ cylinders, the dominant correlation length is consistently charge-neutral, peaking strongly for $\Phi_{\text{ext}}=\pi/2$ external flux threading the cylinder (see inset).
-The maximum extrapolated value of the correlation length increases from $L_{y}=4$ to $6$, consistent with an approach to a continuous transition in the thermodynamic limit.
-It is illustrative to phrase the above discussion in terms of operations on the internal gauge field introduced in the parton construction.
-For simplicity, let us focus on saddle point configurations of the gauge field on each side of the transition, ignoring fluctuations.
-In both the CSL and IQH ground states, the gauge flux through the triangular surface plaquettes should match the external flux, $\phi_{\triangle}=\pi/2$ [6, 7, 12].
-Thus, the only remaining gauge-invariant degree of freedom is the gauge flux through the cylinder rings; assuming $T_{x}^{2}$ is unbroken and $L_{y}$ is odd, we may parameterize the two fluxes as $(\phi,\phi+\pi)$, differing by $\pi$ due to Gauss’s law, as invoked above.
-Since the minimally-entangled CSL ground states are adiabatically connected to the Mott limit in which $\mathcal{P}$ acts trivially, their ring flux patterns are restricted to be $(\pi/2,-\pi/2)$ and $(-\pi/2,\pi/2)$, which are insensitive to the action of $\mathcal{P}$ on ring fluxes, namely $\phi\to\pi-\phi$ (Fig.
+(the spin rotation $i\sigma_{y}$ is included in $\mathcal{P}$ so it leaves the spin operator $\bm{S}=\frac{1}{2}c^{\dagger}\bm{\sigma}c$ invariant; see App.
+C).
+Denoting translation along the two non-parallel directions shown in Fig.
+1 as $T_{x,y}$, which we represent in a gauge where
+$T_{x}c^{\dagger}_{x,y}T_{x}^{-1}=(-1)^{y}c^{\dagger}_{x+1,y},\quad T_{y}c^{\dagger}_{x,y}T_{y}^{-1}=c^{\dagger}_{x,y+1},$ (2)
+we obtain the required magnetic translation algebra $T_{x}T_{y}=(-1)^{N_{F}}T_{y}T_{x}$, where $N_{F}$ is the fermion number.
+The commuting translations $(T_{x})^{2},T_{y}$ define a larger “magnetic” unit cell, yielding two sub-bands with Chern numbers $C=\pm 1$ related by $\mathcal{P}$ (Fig.
+1a).
+We fix the filling to $n=1$ electron per lattice site, corresponding to full filling of the lowest spin-degenerate magnetic sub-band.
+Dimensional reduction can break the $T_{x}$ symmetry.
+Consider the closed “tin can” shaped surface demarcated by the pair of adjacent lattice rings at $x$ and $x+1$.
+Applying Gauss’s law on this surface [34, 35], with $\Phi_{x}$ denoting the external flux through the cylinder ring at $x$, we have $\Phi_{x+1}+2L_{y}\Phi_{\Delta}\equiv\Phi_{x}$ modulo $2\pi$.
+The ring fluxes are therefore uniform when $L_{y}$ is even, but staggered when $L_{y}$ is odd.
+In the latter case, $T_{x}$ is explicitly broken.
+Consider the odd circumference case and fix the ring fluxes to alternate between 0 and $\pi$.
+Though $T_{x}$ is broken, the combined translation-like operation $\mathcal{P}T_{x}$ (“glide-PH”) commutes with the Hamiltonian.
+The IQH state, being adiabatically connected to the unique ground state of the non-interacting Hamiltonian, respects $\mathcal{P}T_{x}$.
+However, the odd-$L_{y}$ CSL phase breaks $\mathcal{P}T_{x}$ spontaneously.
+To see this, note that in the CSL phase the charge fluctuations are gapped; when $U$ is large, the low energy degrees of freedom are described by an effective spin model [6].
+Since $\mathcal{P}\bm{S}\mathcal{P}^{-1}=\bm{S}$, this spin model has a pure translation symmetry $t_{x}$.
+As each ring contains an odd number of spin-$1/2$ moments, the Lieb-Schulz-Mattis theorem requires that any gapped ground state be at least twofold degenerate [36], e.g., by breaking $t_{x}$ spontaneously.
+The same conclusion can be drawn from the anyon content of the CSL: the presence of one semion per unit cell [37, 38, 39, 34] causes $t_{x}$ to permute the two minimally-entangled ground states [40, 41, 42].
+We can also interpret the CSL symmetry breaking in terms of the parton decomposition of the electron into a bosonic chargon $b$ carrying the charge and a fermionic spinon $f$ carrying the spin, $c_{\sigma}=bf_{\sigma}$ [43, 44, 45].
+This requires introducing an internal gauge field $a$.
+In both the CSL and IQH phases, the saddle point configuration of the internal gauge field yields an internal flux through the triangular surface plaquettes that matches the external flux, $\phi_{\triangle}=\pi/2$ [5, 6, 20].
+Moreover, assuming $L_{y}$ is odd, the internal flux through adjacent cylinder rings may be written as $(\phi,\phi+\pi)$, again differing by $\pi$ due to Gauss’s law.
+Since the CSL ground states are adiabatically connected to the Mott limit where $\mathcal{P}$ must act as the identity operator, at strong coupling their internal ring fluxes will energetically lock to $(\pi/2,-\pi/2)$ or $(-\pi/2,\pi/2)$, which are both invariant under the action of $\mathcal{P}$, namely $\phi\to\pi-\phi$ (Fig.
 2a).
-We readily see that these two configurations are related by $\mathcal{P}T_{x}$, exposing the two-fold degeneracy and SSB of the CSL (Fig.
-2b, top) [41].
-In the IQH phase, on the other hand, the Meissner effect pins together the internal and external gauge fields [42].
-The resulting flux pattern $(0,\pi)$ maps into itself under $\mathcal{P}T_{x}$, consistent with a unique, gapped ground state (Fig.
+These two configurations are related by $\mathcal{P}T_{x}\simeq t_{x}$, indicating spontaneous symmetry breaking (SSB) in these odd-$L_{y}$ geometries (Fig.
+2b).
+In the IQH phase where the chargon is condensed, the Higgs mechanism instead pins together the internal and external gauge fields [46, 47, 48].
+The resulting internal ring flux pattern $(0,\pi)$ respects $\mathcal{P}T_{x}$, consistent with a unique, gapped ground state (Fig.
 2c, top).
-While the small $U$ IQH phase does not break the glide-PH symmetry and the intermediate $U$ CSL phase has a two-fold symmetry breaking pattern on odd circumference cylinders, the large $U$ $120^{\circ}$ phase breaks translational symmetry (including the glide PH) in a threefold spatial pattern.
-Numerical observation of phase transition—We first comment on evidence for the phases themselves.
-The primary diagnostic of chiral topological behavior is the behavior of the charge-resolved entanglement spectrum (ES) [43, 44] at a bipartition of the system into two semi-infinite halves, in this case between any pair of adjacent rings [40, 11].
-We exploit the relationship between the physical and entanglement edges [45, 43, 46], namely that the low-lying states in the entanglement spectrum correspond to those of the edge CFT [47, 48, 49, 50].
-Within the bulk-boundary correspondence, the CSL is expected to host gapless edges described at low energies by a free bosonic conformal field theory [51].
-The counting of low-energy states as a function of momentum in each conserved charge sector can be obtained combinatorially: for fixed $n\in\mathbb{Z}_{\geq 0}$ and associated momentum $q_{y}=2\pi n/L_{y}$, one considers a linearly dispersing mode and imagines exciting states from below to above the Fermi level.
-Note that these bosonic objects are charge and spin-neutral but carry momentum.
-The counting sequence is namely the number of integer partitions of $n$: $(1,1,2,3,5,\dots).$ On the other hand, the IQH phase has two independent bosonic edge modes.
-Here, the total momentum must be the sum of those of the two flavors.
-Working out the “two-color” integer partitions of $n$ yields the counting $(1,2,5,10,20\dots).$ These expectations are borne out in our numerics; see Fig.
-2(b,c).
-We now investigate the transition.
-Consistent with the expectations above, we confirm that the CSL phase spontaneously breaks $\mathcal{P}T_{x}$ symmetry on odd-$L_{y}$ cylinders.
-We first find that the bipartite ES in the CSL phase exhibits a two-ring periodicity, which implies SSB.
-In contrast, the ES of the IQH is fully invariant under the action of $\mathcal{P}T_{x}$, as shown in Fig.
-2(c).
-Second, by scanning $U/t$ and inspecting the transfer matrix spectrum, we identify a diverging correlation length for the accessible odd-circumference cylinder geometries, namely $L_{y}=3,5$.
-The divergence is found to belong to the $(Q,S_{z},k_{y})=\bm{0}$ charge sector (see Fig.
-3b for the $L_{y}=5$ and the SM [29] for $L_{y}=3$).
-The spontaneous breaking of glide-PH symmetry is the source of this quasi-(1+1)D criticality.
-To substantiate this, we identified an order parameter $\mathcal{O}_{\delta\mathrm{Heis}}=\sum_{m}\mathcal{O}^{\delta\mathrm{Heis}}_{2m}$, where
+Numerical observation of phase transition—We have used the iDMRG algorithm to determine the ground state on cylinders of circumference $L_{y}=3,4,5,6$.
+For the numerically studied range $U\leq 14$, we find no sign of the strong-coupling $120^{\circ}$ AF, so we focus on the IQH and CSL phases.
+We study the entanglement spectrum (ES) [49, 50] at a spatial bipartition of the system into left and right semi-infinite cylinders [8, 19].
+The low-lying part of the ES corresponds to the conformal field theory (CFT) of the physical edge and exhibits degeneracies characteristic of the edge CFT [51, 50, 52, 53, 54, 55, 56, 57].
+Consistent with the above expectations, the CSL phase spontaneously breaks $\mathcal{P}T_{x}$ symmetry on odd-$L_{y}$ cylinders: the low-lying entanglement eigenstates carry half-integer (integer) spin on bipartitions made to the right of even (odd) cylinder rings $x$ (see App.
+D for details).
+At the even cuts, the number of nearly-degenerate states with fixed $S_{z}=\hbar/2$ are $(1,1,2,3,5,\dots)$ at consecutive momenta; the counting of $S_{z}=0$ states on the odd cuts is similarly $(1,1,2,\dots)$ (see Fig.
+2b, bottom).
+This pattern is characteristic of the free boson CFT [58].
+In contrast, the low-lying ES of the IQH features states with both integer and half-integer spin (App.
+D).
+Consistent with unbroken $\mathcal{P}T_{x}$ symmetry, the spectrum is equivalent on every bipartition (Fig.
+2c, bottom).
+The approximate degeneracies at fixed $S_{z}$ are $(1,2,5,10,20\dots)$, corresponding to two copies of the free boson CFT [59, 60, 61, 62].
+We also construct a $\mathbb{Z}_{2}$ order parameter for $\mathcal{P}T_{x}$, $\mathcal{O}=\sum_{x\in\mathbb{Z}}J^{\perp}_{x}$ where $J^{\perp}_{x}$ is the average current on the circumferential bonds at $x$.
+In the chosen gauge at $\Phi_{\mathrm{ext}}=0$:
+$J^{\perp}_{x}=(-1)^{x}L^{-1}_{y}\sum_{y}i(c_{x,y}^{\dagger}c_{x,y+1}-\mathrm{h.c.}).$ (3)
+Since $(\mathcal{P}T_{x})J^{\perp}_{x}(\mathcal{P}T_{x})^{\dagger}=-J^{\perp}_{x+1}$, then $\mathcal{O}$ is charged under $\mathcal{P}T_{x}$.
+Since the ground states are invariant under $(\mathcal{P}T_{x})^{2}$ so that $\langle J^{\perp}_{x}\rangle=\langle J^{\perp}_{x+2}\rangle$, it suffices to examine the condensation of $\mathcal{O}_{0}=J^{\perp}_{0}+J^{\perp}_{1}$.
+In the CSL phase, $\langle\mathcal{O}_{0}\rangle\neq 0$ (see Fig.
+3a), signaling SSB at $L_{y}=3,5$.
+Near the transition at odd $L_{y}$, the transfer matrix spectrum [63] contains a diverging correlation length in the $(Q,S_{z},k_{y})=0$ charge sector (see Fig.
+3(b) for $L_{y}=5$ and App.
+G for $L_{y}=3$) consistent with the charge of the order parameter $\mathcal{O}$.
+Threading $\Phi_{\text{ext}}=\pi/2$ external flux through the odd-$L_{y}$ cylinder reduces the transition to a crossover since glide-PH symmetry is explicitly broken when $\Phi_{\text{ext}}\not\in\pi\mathbb{Z}$.
+We further measure $\mathcal{O}_{\delta\mathrm{Heis}}=\sum_{m\in\mathbb{Z}}\mathcal{O}^{\delta\mathrm{Heis}}_{2m}$, defined by
 $\mathcal{O}^{\delta\text{Heis}}_{x}=L_{y}^{-1}\sum_{y}\bm{S}_{x,y}\cdot\bm{S}_{x+1,y}-\bm{S}_{x+1,y}\cdot\bm{S}_{x+2,y}.$ (4)
-The operator $\mathcal{O}_{\delta\mathrm{Heis}}$ is both charged under $\mathcal{P}T_{x}$ and condenses in the CSL phase [see Fig.
-3(a)], signaling SSB.
-We remark that threading external flux $\Phi_{\text{ext}}=\pi/2$ through an odd-circumference cylinder reduces the transition to a crossover, owing to the fact that flux threading explicitly breaks glide-PH symmetry unless $\Phi_{\text{ext}}\in\pi\mathbb{Z}$ (see SM [29]).
-On the even-circumference geometries $L_{y}=4$ and $6$, we reproduce the weak crossover, with a correlation length of $\sim 2$ lattice sites, found by Kuhlenkamp et al.
-for $L_{y}=6$ and $\Phi_{\text{ext}}=0$ [11].
-This is consistent with the absence of translation symmetry-breaking in the even-circumference Mott limit.
-Upon threading $\Phi_{\text{ext}}=\pi/2$ external flux through each of these systems, we observe dramatically larger values of the correlation length for $U$ near $U_{c}$ (Fig.
-3), notably in the same neutral sector in which the odd-circumference correlation length diverges [see Fig.
+The order parameter $\mathcal{O}_{\delta\mathrm{Heis}}$ is likewise charged under $\mathcal{P}T_{x}$ and condenses in the CSL phase for odd $L_{y}$ and $\Phi_{\text{ext}}=0$ (see App.
+E).
+The two order parameters share the same symmetries and would be linearly coupled in a Ginzburg-Landau theory.
+We prefer $\mathcal{O}$ defined in Eq.
+(3) because it is larger in magnitude and simpler to express.
+On the even circumference geometries $L_{y}=4,6$ we reproduce the weak crossover, with a correlation length of $\lesssim 1.5$ lattice sites, found by Kuhlenkamp et al.
+for $L_{y}=6$ and $\Phi_{\text{ext}}=0$ [19] (App.
+F1).
+Remarkably, upon threading $\Phi_{\text{ext}}=\pi/2$, the correlation length in the range $10<U<12$ increases dramatically, notably in the same $(Q,S_{z})=0$ sector in which our odd-circumference correlation lengths diverge [see Fig.
 3(c,d)].
-We also perform an extrapolation to infinite bond dimension following Ref.
-[52]; see SM for details [29].
-The maximal extrapolated value is $\xi=6.6$ at $L_{y}=4$, increasing substantially to $\xi=9.1$ at $L_{y}=6$.
-Consistent with the view that accurate estimation of the 2+1D gap and other properties requires consideration of all threaded cylinder fluxes [53], we conjecture that the $\Phi_{\text{ext}}=\pi/2$ threaded system is more reflective of the thermodynamic limit than the unthreaded system, and that larger even-$L_{y}$ geometries will exhibit an increasingly-sharp sequence of crossovers toward 2+1D critical behavior in the charge-neutral sector.
-Figure 4: Density-density [left panel] and spin-spin [right panel] connected correlations for odd-circumference cylinders with $U$ tuned to the respective critical points, $U_{*}(L_{y}=3)=11.1$ and $U_{*}(L_{y}=5)=11.6$.
-Both $L_{y}=5$ correlation functions are scaled down by 100 for ease of viewing.
-(a) Density-density correlations decay nearly algebraically at the IQH-CSL transition, pointing to the presence of gapless charge-neutral excitations in the 2+1D limit.
-(b) In contrast, correlations of the spin density decay exponentially with no apparent enhancement from $L_{y}=3$ to $5$, suggesting that the excitations are spin-rotationally-invariant singlets.
-Remarkably, the critical fluctuations at the odd-$L_{y}$ critical points manifest not only in the order parameter, but also in the charge density.
-This is seen manifestly in the connected density-density correlations $\langle n(x)n(0)\rangle$, with $n(x)$ averaged over the ring $x$, which we find to decay approximately algebraically, i.e., linearly on a log-log scale [see Fig.
-4(a)] However, estimation of the 2+1D critical exponents from these data is hampered by the fact that the critical features of the thermodynamic limit need only show up in the regime $1\ll x\ll L_{y}$, which is inaccessible for the presently-tractable system sizes.
-While PH-glide is no longer broken for $L_{y}=4,6$ and the operator Eq.
-(3) no longer condenses (see Fig.
-3a), we also observe that their enhanced correlation length is accompanied by an apparent algebraic plateau in $\langle n(x)n(0)\rangle$ over the scale of the correlation length (see SM [29]).
-Curiously, we do not observe critical behavior in the odd-$L_{y}$ connected correlations of the spin density $S_{z}$ [Fig.
-4(b)], which is also neutral under the conserved charge and spin $U(1)$ symmetries.
-However, since $S_{z}$ may be rotated into $S_{x}$ by a global spin rotation, under which the CSL is invariant, this is consistent with our not observing a diverging correlation length in the $(Q,S_{z})=(0,\hbar)$ sector of the transfer matrix.
-Altogether, under the assumption that this behavior persists for the entire sequence of odd circumferences, our observations suggest that the 2+1D critical point hosts gapless, charge-neutral spin-singlet excitations.
-We remark that these data are consistent with the partonic view that the transition into the IQH corresponds to chargon condensation.
-While the chargon is not gauge invariant and may not develop conventional long-range connected correlations, at the critical point the combination of the chargon, anti-chargon and gauge string is a gapless neutral excitation [42].
-Interestingly, the power law correlations presented in Fig.
-4 show that the physical charge density operator, which has nonzero overlap with the gauge-invariant chargon density $b^{\dagger}(x)b(x)$, produces a state with a nonzero overlap with the critical excitations when applied to the ground state, raising the intriguing possibility that the critical fluctuations can be experimentally observed.
-Discussion—By combining numerical data on both even and odd-circumference cylinder geometries, we have provided evidence that the CSL-IQH transition is in fact continuous in the Hofstadter-Hubbard model at $\pi/2$ flux per triangle and that critical properties are already manifest at the numerically-accessible circumferences $3\leq L_{y}\leq 6$.
-Our finding that critical fluctuations appear in the charge density correlation function, which may be probed by experimental or numerical measurement of the structure factor, raise the possibility of experimental and numerical investigation of this transition.
-This paradigmatic Hofstadter-Hubbard system provides a uniquely minimal and experimentally-relevant setting [10, 11] in which to study a fundamental phase transition at the interface of spin liquid and quantum Hall physics.
-As such, a more complete understanding of this transition is an important task for future work.
-Quantitatively, our investigation clarifies the possible observable, numerical signatures that may be used to probe the phase transition in a variety of numerical methods, such as AFQMC, VMC, or fermionic PEPS, which may be able to access larger system sizes or spectral signatures of spin-charge separation [54, 55, 56].
-An important broader message of this work is that the ability to extrapolate 2+1D physics from dimensionally-reduced numerical data—namely on infinite cylinder geometries—is greatly enhanced by carefully considering the manner in which candidate ground states and their phase transitions respond to dimensional reduction and to external fields that become boundary effects in the thermodynamic limit.
-Applying such techniques more systematically, it would be valuable to characterize the contexts in which UV features of Mott phases, namely spin liquids, may be leveraged to better characterize their charge fluctuation-induced phase transitions.
+At $L_{y}=4$, extrapolation to infinite bond dimension following Ref.
+[64] predicts a maximal extrapolated value of $\xi_{L_{y}=4}^{\chi\to\infty}=6.6$ in units of cylinder rings (see App.
+F2).
+While at $L_{y}=6$, $\xi=5.8$ at the largest bond dimension $\chi=32768$, the correlation length is still growing rapidly with $\chi$.
+This hampers the precise extrapolation of $\xi_{L_{y}=6}^{\chi\to\infty}$ but also indicates it substantially exceeds $\xi_{L_{y}=4}^{\chi\to\infty}$.
+This suggests that the larger even-$L_{y}$ geometries will exhibit a sequence of increasingly-sharp crossovers toward 2+1D criticality.
+We now discuss the behavior of correlation functions.
+The connected current correlator $C_{JJ}(x)=\langle J_{0}^{\perp}J_{x}^{\perp}\rangle_{\text{c}}$ exhibits power-law decay at the approximate transition points $U_{c}(L_{y}=3)=11.1$ and $U_{c}(L_{y}=5)=11.6$ (Fig.
+4a), chosen as $U=U_{c}$ where the correlation length is largest at the largest available $\chi$.
+The critical fluctuations also manifest in the charge density: the connected correlations $C_{nn}(x)=\langle n_{0}n_{x}\rangle_{\text{c}}$ (with $n_{x}$ averaged over ring $x$) are negative and decay algebraically (Fig.
+4b), but fall off much more rapidly than the current correlations.
+While there is no SSB for $L_{y}=4,6$ we still observe an enhanced correlation length (see Fig.
+3) and slow decay of the current and density correlations (App.
+H).
+On the other hand, the connected spin correlations $C_{S^{z}S^{z}}(x)=\langle S^{z}_{0}S^{z}_{x}\rangle_{\text{c}}$ (with $S^{z}_{x}$ averaged over ring $x$) exhibit clear exponential decay at the odd-$L_{y}$ critical points, oscillating in sign with a period of two rings, with no apparent enhancement from $L_{y}=3$ to $L_{y}=5$ (Fig.
+4c).
+The electron correlators $C_{c^{\dagger}c}(x)=L_{y}^{-2}\sum_{k_{y}}|\langle c^{\dagger}(0,k_{y})c(x,k_{y})\rangle|$ likewise decay exponentially (Fig.
+4d), so that both spin and electron excitations are gapped at the transition.
+Should these behaviors persist in the thermodynamic limit, they imply that the 2+1D critical point is continuous and hosts gapless, charge-neutral spin-singlet fluctuations, namely of the current.
+Figure 4: Connected correlation functions on the odd $L_{y}$ cylinders at their critical point, $U_{c}(L_{y}=3)=11.1$ (yellow) and $U_{c}(L_{y}=5)=11.6$ (blue).
+The $L_{y}=5$ correlations are reduced by $\times 10$ for ease of viewing.
+Upper: Current-current correlations and negative of density-density correlations, approaching quasi-long range order as a function of MPS bond dimension $\chi$ (increasing light to dark).
+Lower: Spin-spin (plotted with alternating sign $(-1)^{x}$) and electron-hole correlations decay exponentially, demonstrating a spin gap.
+We comment briefly on the nature of the odd-$L_{y}$ quasi-(1+1)(1+1)D critical points.
+As the $1e$ and $2e$ correlators remain gapped (App.
+G1), the data are incompatible with Luttinger and Luther-Emery liquid criticality [65, 66].
+The spontaneous breaking of a $\mathbb{Z}_{2}$ symmetry suggests the transition is in the 2D Ising universality class.
+The “ionic” Hubbard chain [67, 68, 69] shares symmetries and behavior with the studied odd $L_{y}$ systems [70, 71, 72, 73] and has a Mott transition; while the model has subtleties that have complicated the unambiguous determination of exponents, it is likely that the transition is of the Ising type (see App.
+G2 for a more detailed discussion).
+Further investigation into the odd $L_{y}$ sequence of transitions, particularly precisely how they evolve to the 2+1D transition and which critical correlations survive, is needed.
+The prominence of current-current correlations, namely their slow falloff at both short and long distances, suggests they remain prominent at the 2+1D critical point.
+Discussion—Our numerical data on both even and odd-circumference cylinder geometries provides confirming evidence of a CSL phase and indicates that the CSL-IQH transition is in fact continuous in the Hofstadter-Hubbard model at $\pi/2$ flux per triangle.
+The single-electron and spin gaps remain nonzero across the transition, while the current-current correlation function exhibits power-law behavior at the odd-$L_{y}$ critical points and is also prominent on the even-$L_{y}$ geometries.
+This raises the possibility of experimental detection of the transition and provides valuable guidance for its further numerical examination.
+Using quantum Monte Carlo, fermionic PEPS, or neural quantum states, one may be able to access wider systems, yielding confirming evidence of quantum criticality, estimates of critical exponents, and possible signatures of conformal invariance.
+A broader message of this work is that the ability to extrapolate 2+1D physics from dimensionally-reduced data is greatly enhanced by analyzing how symmetries, ground states and their phase transitions respond to spatial compactification and flux threading.
+In particular, the properties of transitions into topologically-ordered phases may generally be illuminated by studying cylinder compactifications in which topological ground state degeneracy manifests as SSB [74, 75].
+Leveraging such techniques systematically, it would be valuable to characterize the contexts in which UV features of Mott and quantum Hall phases offer improved understanding of their charge fluctuation-driven quantum phase transitions.
+Acknowledgements.
 Acknowledgements—We acknowledge useful discussions with E.
@@
 Chatterjee, O.
-Gauthé, M.
+Gauthé, Y-C.
+He, M.
 Knap, C.
@@
 Liu, F.
-Pollmann, Z.
-Weinstein, and especially X.Y.
-Song.
-The Flatiron Institute is a division of the Simons Foundation.
-AJM was supported in part by Programmable Quantum Materials, an Energy Frontier Research Center funded by the U.S.
-Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0019443.
+Pollmann, R.
+Sahay, R.
+Verresen, Z.
+Weinstein, Y-H.
+Zhang, and especially X-Y.
+Song and A.
+Vishwanath.
 S.D.
 acknowledges support from the NSERC PGSD fellowship.
+A.J.M.
+was supported in part by Programmable Quantum Materials, an Energy Frontier Research Center funded by the U.S.
+Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0019443.
 M.Z.
@@
 This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8683 to T.S.
-Calculations were performed using the TeNPy Library [57].
+Calculations were performed using the TeNPy Library [76, 77, 78].
+The Flatiron Institute is a division of the Simons Foundation.
 Supplementary Material: “Chiral Spin Liquid and Quantum Phase Transition in the Triangular Lattice Hofstadter-Hubbard Model”
 Appendix A Hamiltonian construction at all rational fluxes
-For the underlying triangular lattice — with the “YC” orientation on the cylinder (where the $\hat{y}$ site-to-site distance is the NN distance $a$) — we take the following Bravais vectors:
-[…]
-whose corresponding reciprocal lattice vectors are
-[…]
-SI Choosing the magnetic unit cell
-The following discussion applies to both the YC and XC lattice geometries. Let the flux per parallelogram be
+SI YC Lattice Parameterization
+We first describe two coordinate parameterizations of the underlying triangular lattice, beginning in the 2D limit.
+In the “YC” parameterization of the 2D plane, the $\hat{y}$ site-to-site distance is the nearest-neighbour distance $a$.
+This corresponds to the Bravais vectors
+$\bm{a}_{1}=a\left(\frac{\sqrt{3}}{2}\hat{x}+\frac{1}{2}\hat{y}\right),\qquad\bm{a}_{2}=a\hat{y}\qquad\text{(YC)},$ (1)
+and the corresponding reciprocal lattice vectors
+$\bm{b}_{1}=\frac{4\pi}{\sqrt{3}a}\hat{x},\qquad\bm{b}_{2}=\frac{2\pi}{a}\left(-\frac{\hat{x}}{\sqrt{3}}+\hat{y}\right)\qquad\text{(YC)}.$ (2)
+This leads to a natural cylinder compactification: one may identify each point
+$\bm{r}=x\bm{a}_{1}+y\bm{a}_{2}$ (3)
+with points $\bm{r}^{\prime}$ for which $y\equiv y^{\prime}\ (\text{mod }L_{y})$ and also $x=x^{\prime}$.
+We refer to $L_{y}$ as the “circumference” of the cylinder.
+Doing so for the YC lattice, one obtains the canonical “YC-$L_{y}$” cylinder geometry [8], whose circumference is $L_{y}a$.
+SII Choosing the magnetic unit cell
+Let the flux per parallelogram (i.e., two triangles) be
 $\Phi=\frac{p}{q}\Phi_{0},$ (4)
@@
 $B=\frac{\Phi}{A_{\text{par}}}=\frac{p\Phi_{0}/q}{\frac{\sqrt{3}}{2}a^{2}}=\frac{2p\Phi_{0}}{\sqrt{3}q}a^{-2}.$ (5)
-Regardless of the parity of $q$, there will be $q$ magnetic sublattice sites in the magnetic unit cell, which we label by $\tau=0,\dots,q-1.$ When $q$ is even, the magnetic Bravais vectors are
+Regardless of the parity of $q$, there will be $q$ magnetic sublattice sites in the magnetic unit cell. When $q$ is even, the magnetic Bravais vectors are
 $\bm{R}_{1}=q\bm{a}_{1}-\frac{q}{2}\bm{a}_{2}\propto\hat{x},\qquad\bm{R}_{2}=\bm{a}_{2}\propto\hat{y},\qquad(q\ \text{even}),$ (6)
-whose total area $A_{\text{muc}}$ (for either YC or XC) is precisely that of $2q$ triangles of area $\frac{\sqrt{3}}{4}a^{2}.$ The associated reciprocal lattice vectors are
+whose total area $A_{\text{muc}}$ is precisely that of $2q$ triangles of area $\frac{\sqrt{3}}{4}a^{2}.$ The associated reciprocal lattice vectors are
 $\bm{G}_{1}=\frac{1}{q}\bm{b}_{1}\propto\hat{x},\qquad\bm{G}_{2}=\frac{1}{2}\bm{b}_{1}+\bm{b}_{2}\propto\hat{y},\qquad(q\ \text{even}).$ (7)
@@
 When $q$ is odd, the enclosing magnetic unit cell again contains exactly $2q$ triangles, but we take
-$\bm{R}_{1}=q\bm{a}_{1}-\frac{q-1}{2}\bm{a}_{2},\qquad\bm{R}_{2}=\bm{a}_{2}\propto\hat{y},,\qquad(q\ \text{odd}),$ (8)
+$\bm{R}_{1}=q\bm{a}_{1}-\frac{q-1}{2}\bm{a}_{2},\qquad\bm{R}_{2}=\bm{a}_{2}\propto\hat{y},\qquad(q\ \text{odd}),$ (9)
 whose total area is once again
@@
 $\bm{G}_{1}=\frac{1}{q}\bm{b}_{1},\qquad\bm{G}_{2}=\frac{q-1}{2q}\bm{b}_{1}+\bm{b}_{2},\qquad(q\ \text{odd}),$ (11)
-which also satisfy the inverse area relationship, Eq. (7).
-In Fig.
-LABEL:fig:YC_c2zsymm_locations, we illustrate the choice of magnetic unit cell for both odd and even $q$ for the YC triangular lattice boundary conditions.
+which also satisfy the inverse area relationship, Eq.
+(8).
 For both even and odd denominator $q,$ note that the enclosing magnetic unit cell contains exactly $2q$ triangles, so $2\pi p\in 2\pi\mathbb{Z}$ total flux.
-In addition, we position the sublattice sites $\tau$ within the magnetic unit cell in such a way that all the sites within magnetic unit cell $\bm{R}$ transform under $C_{2z}$ into sites in $-\bm{R}$.
-In particular, the sites in the home magnetic unit cell transform into one another in a clean way.
-We achieved this by further partitioning the odd-$q$ geometries into two subclasses.
-In Fig.
-LABEL:fig:XC_c2zsymm_locations, we make the same specifications for the XC boundary conditions.
-SII Hamiltonian and choice of vector potential
-Placing the underlying triangular lattice on the cylinder, let the resulting system have circumference $L_{y}$, i.e., the points $\bm{r}\sim\bm{r}+L_{y}\bm{a}_{2}$ are geometrically identified.
-Let the lattice sites be indexed by
-$\bm{r}=x\bm{a}_{1}+y\bm{a}_{2},$ (11)
-which defines our coordinate system, where $(x,y)\in\mathbb{Z}^{2}.$ Since $\bm{a}_{2}$ is purely vertical in both geometries, then $x$ indexes the rings, and $y$ is a integer spatial index within each ring.
-To generate complex hoppings and modify the translation algebra to a magnetic translation algebra, we introduce an external $U(1)$ vector potential.
+SIII Hamiltonian and choice of vector potential
+In all cases, we adopt the standard convention for the on-site Hubbard repulsion:
+$H_{\mathrm{int}}=U\sum_{\bm{r}}n_{\bm{r}\uparrow}n_{\bm{r}\downarrow},\qquad U\geq 0.$ (12)
+Now we discuss the hoppings.
+Placing the underlying triangular lattice on the cylinder as described in Sec.
+SI, let the resulting system have circumference $L_{y}$, i.e., the points $\bm{r}\sim\bm{r}+L_{y}\bm{a}_{2}$ are geometrically identified.
+Let the lattice sites be indexed as in Eq.
+(3), which defines a set of coordinate coefficients $(x,y)\in\mathbb{Z}^{2}.$ Since $\bm{a}_{2}$ is purely vertical, then $x$ indexes the cylinder “rings”, and $y$ is an integer-valued spatial index within each ring.
+To generate complex hoppings and modify the translation algebra into a magnetic translation algebra, we introduce an external $U(1)$ vector potential.
 We assume that it is linear and that its curl is the magnetic field:
@@
 where $B$ is as in Eq. (5). The hoppings are obtained from the vector potential via the Peierl’s substitution, for which we adopt the following sign convention:
-[…]
+$t_{\bm{r}^{\prime}\leftarrow\bm{r}}=t\exp\left(i\int^{\bm{r}^{\prime}}_{\bm{r}}\bm{A}(\bm{\ell})\cdot d\bm{\ell}\right)=t\exp\left(i\bm{A}\left(\frac{\bm{r}^{\prime}+\bm{r}}{2}\right)\cdot(\bm{r}^{\prime}-\bm{r})\right),$ (14)
 where the second equality is due to linearity of $\bm{A}(\bm{r})$. Our convention is that the single-particle Hamiltonian is (note the minus sign):
@@
 $t_{\bm{r}^{\prime}\leftarrow\bm{r}}=te^{i\bm{A}\left(\frac{\bm{r}^{\prime}+\bm{r}}{2}\right)\cdot(\bm{r}^{\prime}-\bm{r})}.$ (16)
-To constrain the form Eq.
-(12), we remark that the terms $x\hat{x}$ and $y\hat{y}$ are curl-free, whereas $Bx\hat{y}$ and $-By\hat{x}$ — the standard Landau gauges on the rectangular lattice — have the correct curl.
-We therefore reason that the most general permissible $M$ is given by
+To constrain the form of the vector potential, we remark that the terms $x\hat{x}$ and $y\hat{y}$ are curl-free, whereas $Bx\hat{y}$ and $By\hat{x}$ (the standard Landau gauge potentials on the rectangular lattice) have the correct curl.
+We therefore reason that the most general permissible $M$ in Eq.
+(13) is given by
 $M=B\begin{pmatrix}\alpha&\gamma-1\\ \gamma&\beta\end{pmatrix}$, (17)
@@
 We will now demand that the choice of vector potential is compatible with our chosen geometry of magnetic Bravais vectors.
-Specifically, we ask that the Peierl’s hoppings generated by $\bm{A}$ are periodic in the chosen Bravais vectors.
-This allows us to place the system easily on a torus or cylinder, provided the system tiles an integer number of magnetic unit cells.
+Specifically, we ask that the Peierl’s hoppings generated by $\bm{A}$ are periodic in the chosen magnetic Bravais vectors.
+This allows us to place the system easily on a torus or cylinder, provided the system is an integer tiling of magnetic unit cells.
 Quantitatively, we demand that the RHS of Eq.
@@
 [79], an elegant solution can be found in terms of the reciprocal vectors of the underlying triangular lattice.
-We take the following since $\bm{b}_{1}\propto\hat{x}$ for both YC and XC geometries, and we want the vector potential to be independent of $y$:
-$\bm{A}(\bm{r})=\frac{\Phi}{(2\pi)^{2}}(\bm{r}\cdot\bm{b}_{1})\bm{b}_{2}.$ (18)
+We take the following since $\bm{b}_{1}\propto\hat{x}$, and we want the vector potential to be independent of $y$:
+$\bm{A}(\bm{r})=\frac{\Phi}{(2\pi)^{2}}(\bm{r}\cdot\bm{b}_{1})\bm{b}_{2},$ (19)
 where $\Phi$ is the flux per parallelogram. Noting the vector identity
-[×$\left[\nabla\times\left((\bm{r}\cdot\bm{Q})\bm{P}\right)\right]_{i}$ $=\epsilon_{ijk}\partial_{j}\left(\bm{r}_{\ell}\bm{Q}_{\ell}\bm{P}_{k}\right)=\epsilon_{ijk}\bm{Q}_{j}\bm{P}_{k}=(\bm{Q}\times\bm{P})_{i},$ (19)
+$\left[\nabla\times\left((\bm{r}\cdot\bm{Q})\bm{P}\right)\right]_{i}$ $=\epsilon_{ijk}\partial_{j}\left(\bm{r}_{\ell}\bm{Q}_{\ell}\bm{P}_{k}\right)=\epsilon_{ijk}\bm{Q}_{j}\bm{P}_{k}=(\bm{Q}\times\bm{P})_{i},$ (20)
 then
-$\bm{B}(\bm{r})$ =×$=\nabla\times\bm{A}(\bm{r})=\frac{\Phi}{(2\pi)^{2}}\bm{b}_{1}\times\bm{b}_{2}=\frac{\Phi}{(2\pi)^{2}}\frac{(2\pi)^{2}}{A_{\text{par}}}\hat{z}=\frac{\Phi}{A_{\text{par}}}\hat{z}.$ (20)
-Moreover, it can be shown to have the required periodicity property:
+$\bm{B}(\bm{r})$ $=\nabla\times\bm{A}(\bm{r})=\frac{\Phi}{(2\pi)^{2}}\bm{b}_{1}\times\bm{b}_{2}=\frac{\Phi}{(2\pi)^{2}}\frac{(2\pi)^{2}}{A_{\text{par}}}\hat{z}=\frac{\Phi}{A_{\text{par}}}\hat{z}$.
+(21)
+Moreover, it can be shown to have the required periodicity property. Note that
 $\bm{a}_{i}\cdot\bm{A}(\bm{R}_{j})$ $=\frac{\Phi}{(2\pi)^{2}}(\bm{R}_{j}\cdot\bm{b}_{1})\bm{b}_{2}\cdot\bm{a}_{i}=\delta_{i,2}\frac{2\pi p/q}{2\pi}(\bm{R}_{j}\cdot\bm{b}_{1})=\delta_{i,2}\frac{p}{q}\bm{R}_{j}\cdot\bm{b}_{1},$ (22)
@@
 $\bm{a}_{i}\cdot\bm{A}(\bm{R}_{j})\in 2\pi\mathbb{Z}.$ (24)
+Specializing to the case $(p,q)=(1,2)$ relevant to the main text, we plot the resulting hopping amplitudes Eq.
+(16) in Fig.
+S1(a) for the YC parameterization of the 2D lattice.
+The hoppings only take values in $\{\pm 1,\pm i\}$ and are evidently periodic in both $\bm{R}_{1}$ and $\bm{R}_{2}$.
+The color of each bond denotes the complex phase of $t_{\bm{r}^{\prime}\leftarrow\bm{r}}$ (see Eq.
+(15)) in the direction of the arrow (pictured at the bond centers).
+Figure S1: (a) Graphical depiction of the complex phases of the nearest-neighbor hopping amplitudes of Eq.
+(16) for the $(p,q)=(1,2)$ two-band Hamiltonian in a portion of the infinite 2D plane, where the orientation of the plane is consistent with the YC coordinate system specified by Eqs.
+(1,6).
+The color of each bond denotes the complex phase of $t_{\bm{r}^{\prime}\leftarrow\bm{r}}$ (see Eq.
+(15)) in the direction of the arrow (pictured at the bond centers).
+The two orange dots indicate the two sites in the magnetic unit cell.
+Note that the top/bottom and right/left edges are not implicitly identified.
+(c) Graphical depiction of the fluxes through adjacent loops of a YC-$L_{y}$ cylinder, defined in Sec.
+SI.
+The outside surface of each cylindrical segment has a total area of $2L_{y}$ triangles, each having $\pi/2$ flux [pictured].
+When $L_{y}$ is odd [upper scenario], this leads to a $\pi$-staggering of the flux through each cylinder ring due to Gauss’ law.
+When $L_{y}$ is even [lower scenario], the flux through each ring is the same.
 Appendix B Threading flux
@@
 When dimensionally-reducing the lattice system to the cylinder, the charge and spin flux through the cross sections of the cylinder distinguish gauge-inequivalent systems.
-This additional flux can be implemented by adding a constant term to the vector potential:
+This additional flux can be implemented by adding a constant offset to the vector potential:
 $\bm{A}(\bm{r})\longrightarrow\bm{A}_{\phi}(\bm{r})=\bm{A}(\bm{r})+\phi\frac{\bm{b}_{2}}{2\pi L_{y}},$ (25)
@@
 SII Twisted boundary condition implementation
-For the purposes of numerical implementation, particularly in hybrid space where we exchange $y$ for $k_{y}$, it is convenient to instead implement flux threading via a twist in the boundary conditions.
-Starting with periodic boundary conditions $c^{\dagger}(\bm{r}+L_{y}\bm{a}_{2})=c^{\dagger}(\bm{r})$ for $\phi=0$, define a new set of operators
+For the purposes of numerical implementation, particularly in hybrid space where we trade $y$ for $k_{y}$, it is convenient to instead implement flux threading via a twist in the boundary conditions.
+Starting with periodic boundary conditions for the fermion operators at $\phi=0$, namely $c^{\dagger}(\bm{r}+L_{y}\bm{a}_{2})=c^{\dagger}(\bm{r})$, we define a new set of operators
 $c^{\dagger}_{\phi}(\bm{r})=c^{\dagger}(\bm{r})\exp(iy\phi/L_{y})=c^{\dagger}(\bm{r})\exp\left(i\phi\frac{\bm{b}_{2}\cdot\bm{r}}{2\pi L_{y}}\right).$ (30)
 Then instead of implementing Eq.
-(24), we can instead replace every instance of $c$ with $c_{\phi}$.
+(25), we can instead replace every instance of $c$ in the second-quantized interacting Hamiltonian with $c_{\phi}$.
 This does not affect terms that depend solely on charge and spin densities, e.g., the Hubbard interaction.
-However, it will modify the hoppings in a way that is equivalent to Eq.
+However, it modifies the hoppings in a way that is equivalent to Eq.
 (25).
@@
 The utility of this alternative formulation is that it leads to a clean numerical implementation upon fourier transforming $y\to k_{y}$.
-In particular, recall that $(x,y)=(\bm{r}\cdot\bm{b}_{1}/2\pi,\bm{r}\cdot\bm{b}_{2}/2\pi)$ and define
+In particular, note that $(x,y)=(\bm{r}\cdot\bm{b}_{1}/2\pi,\bm{r}\cdot\bm{b}_{2}/2\pi)$ and define
 $c_{\phi}^{\dagger}(x,k_{y})=\frac{1}{\sqrt{L_{y}}}\sum_{y}e^{-ik_{y}y}c_{\phi}^{\dagger}(x,y)\iff c_{\phi}^{\dagger}(x,y)=\frac{1}{\sqrt{L_{y}}}\sum_{k_{y}}e^{ik_{y}y}c_{\phi}^{\dagger}(x,k_{y}).$ (35)
 Note that this is the Fourier convention of TeNPy.
-When we express the Hamiltonian in the hybrid $xk_{y}$ representation, we find that all that’s required is to shift the allowed values of $k_{y}.$ To see this, we impose the twisted boundary condition Eq.
+When we express the Hamiltonian in the hybrid $xk_{y}$ representation, we find that all that is required is to shift the allowed values of $k_{y}.$ To see this, we impose the twisted boundary condition Eq.
 (34) on the Fourier-expanded operators:
@@
 $e^{ik_{y}L_{y}}=e^{i\phi}.$ (37)
-The allowed values of momenta are therefore quantized to:
+The allowed values of momenta are therefore quantized to
 $k_{y}=\frac{2\pi}{L_{y}}\left(n+\frac{\phi_{q}}{2\pi}+\frac{\sigma_{z}\phi_{s}}{4\pi}\right),\qquad n\in\mathbb{Z},$ (38)
 which are shifted relative to the $\phi=0$ case.
-When the original Hamiltonian $\hat{H}_{\phi=0}$ is written in hybrid space, some coefficients will depend on $k.$ The only affect of the replacement Eq.
-(29) is then for those coefficients to instead be evaluated on these shifted set of allowed values.
+When the original Hamiltonian $\hat{H}_{\phi=0}$ is written in hybrid space, some coefficients will depend on $k.$ The only effect of the replacement Eq.
+(30) is then for those coefficients to instead be evaluated on this shifted set of allowed values [8].
 Appendix C Two-band hopping Hamiltonian
+Figure S2: (a) Plot of the lower and upper energy bands of the Hofstadter system with $(p,q)=(1,2)$, corresponding to the case with $\Phi_{\triangle}=\pi/2$ considered throughout the main text.
+Momenta are given within the first “magnetic” Brillouin zone, defined with respect to the magnetic reciprocal vectors of Eq.
+(7) in the 2D planar YC coordinate system of Eq.
+(2).
+(b) Plot of the Berry curvature of the aforementioned bands which, like their energy, is equal and opposite at every momentum.
+The bands carry opposite Chern numbers $\pm 1$.
 SI Hamiltonian at $(p,q)=(1,2)$
-Now let us specialize to the case $(p,q)=(1,2)$, which corresponds to $\Phi_{\triangle}=\pi/2$.
-To specify the unit cells (which each contain two sites), we often write $x=\tau+2u$, where $u\in\mathbb{Z}$, while $\tau\in\{0,1\}$ specifies the even/odd rings.
-With periodic boundary conditions and before flux insertion, the hoppings Eq.
+Now let us specialize to the YC lattice geometry and to the case $(p,q)=(1,2)$, which corresponds to $\Phi_{\triangle}=\pi/2$.
+To specify the unit cells (which each contain two sites), we write $x=\tau+2u$, where $u\in\mathbb{Z}$, while $\tau\in\{0,1\}$ specifies the even/odd rings.
+With periodic boundary conditions and absent any flux insertion, the hoppings Eq.
 (16) take the particular form
@@
 $=t\exp\left(-i\bm{A}\left(\frac{\bm{a}_{1}-\bm{a}_{2}}{2}+x\bm{a}_{1}+y\bm{a}_{2}\right)\cdot\bm{a}_{2}\right)$
-$=t\exp\left(-i\pi(x+1/2)\right).$ (40)
-We depict the resulting hopping phases in Fig.
-S1(a) for the lattice orientation where $\bm{a}_{2}=a\hat{y}$ is a nearest-neighbour bond; this becomes becomes the YC geometry when the system is placed on a cylinder under the identification $\bm{r}\sim\bm{r}+L_{y}\bm{a}_{2}$.
-Figure S1: (a) Hopping phases for the $(p,q)=(1,2)$ two-band Hamiltonian corresponding to $\pi/2$ flux per triangle.
-The two black circles indicate the two sites in the magnetic unit cell.
-The phases indicated by the color bar are accrued when hopping upward, right-upward, and right-downward; the complex-conjugate phases are accrued when hopping against these directions.
-(b) Graphical depiction of the fluxes through various plaquettes upon placing the lattice on a cylinder.
-Every triangle on the surface of the infinite cylinder is penetrated by $\Phi_{\triangle}=\pi/2$ external flux.
-As a consequence of Gauss’s law, the external flux through adjacent rings—each of which is a polygonal surface bounded by a ring of length $L_{y}$—must alternate by $\pi$ when $L_{y}$ is odd, and must be uniform when $L_{y}$ is even.
+$=t\exp\left(-i\pi(x+1/2)\right),$ (41)
+as shown graphically in Fig.
+S1(a).
+We depict the resulting energy bands $E_{\text{lower/upper}}(\bm{k})$ in Fig.
+S2, as well as their corresponding Berry curvature $\Omega_{\text{lower/upper}}(\bm{k})$, in the YC coordinate system consistent with Eqs.
+(2) and (7).
+The energies are equal and opposite at every momentum and are gapped everywhere in the Brillouin zone.
+The Berry curvature is also everywhere equal and opposite, with the Chern numbers evaluating to +1+1 for the lower band and 1-1 for the upper band.
 SII Symmetries in the 2D plane
@@
 $=\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{\delta}\right)t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}.$ (48)
-This affects the hopping directions with some weight along $\bm{\delta}=\bm{a}_{2}$ and, by comparison to Eq.
+This affects the hoppings in directions with some weight along $\bm{\delta}=\bm{a}_{2}$ and, by comparison to Eq.
 (29) above, corresponds to threading $\phi=\pi L_{y}$ through the cylinder.
-In the 2D plane geometry — and later on some even-$L_{y}$ geometries as we’ll show — the Hamiltonian can be restored by composing the bare $T_{x}$ with a large gauge transformation.
+In the 2D plane geometry—and on cylinder geometries of even circumference as we’ll later show—the Hamiltonian can be restored by composing the bare $T_{x}$ with a large gauge transformation.
 In particular, define the magnetic translation operation by
-$T_{x}^{\text{mag}}c_{\bm{r}}^{\dagger}(T_{x}^{\text{mag}})^{-1}=c_{\bm{r}+\bm{a}_{1}}^{\dagger}\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right)=c_{\bm{r}+\bm{a}_{1}}^{\dagger}(-1)^{y},$ (48)
+$T_{x}c_{\bm{r}}^{\dagger}T_{x}^{-1}=c_{\bm{r}+\bm{a}_{1}}^{\dagger}\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right)=c_{\bm{r}+\bm{a}_{1}}^{\dagger}(-1)^{y},$ (49)
 so that
-$T_{x}^{\text{mag}}\hat{h}(T_{x}^{\text{mag}})^{-1}$ $=T_{x}^{\text{mag}}\left(\sum c_{\bm{r}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\bm{r}}\right)(T_{x}^{\text{mag}})^{-1}$ (49)
+$T_{x}\hat{h}T_{x}^{-1}$ $=T_{x}\left(\sum c_{\bm{r}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\bm{r}}\right)T_{x}^{-1}$ (50)
 $=\sum\exp\left(\frac{i}{2}\bm{b}_{2}\cdot(\bm{r}+\bm{\delta})\right)c_{\bm{r}+\bm{a}_{1}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\bm{r}+\bm{a}_{1}}\exp\left(-\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right)$ (51)
@@
 SII.2 Particle-hole
-Consider the naive definition of the particle-hole operation:
+Consider the most naive definition of the particle-hole operation:
 $\tilde{\mathcal{P}}c^{\dagger}(\bm{r})\tilde{\mathcal{P}}^{-1}=c(\bm{r}),\qquad\tilde{\mathcal{P}}i\tilde{\mathcal{P}}^{-1}=+i,$ (55)
-where the condition specifies it is a linear (as opposed to anti-linear) unitary many-body operator. If 1 and 2 are an adjacent pair of sites, then particle-hole acts as
+where the second condition specifies that it is a linear (as opposed to anti-linear) unitary many-body operator.
+If 1 and 2 denote an adjacent pair of sites, then this naive particle-hole operation acts as
 $\tilde{\mathcal{\mathcal{P}}}(t_{12}c_{1}^{\dagger}c_{2}+t_{12}^{*}c_{2}^{\dagger}c_{1})\tilde{\mathcal{\mathcal{P}}}^{-1}$ $=t_{12}c_{1}c_{2}^{\dagger}+t_{12}^{*}c_{2}c_{1}^{\dagger}$ (56)
@@
 $=-t_{12}^{*}c_{1}^{\dagger}c_{2}-t_{12}c_{2}^{\dagger}c_{1},$ (58)
-so that it flips the Peierl’s phase and increases it by $\pi.$ On the triangular lattice, this modifies the flux through each triangle as follows:
+so that it reverses the Peierl’s phase and shifts it by $\pi.$ On the triangular lattice, this modifies the flux through each elementary triangular plaquette as follows:
 $\Phi_{\triangle}\to-\Phi_{\triangle}+\pi.$ (59)
-Requiring this operation to be invariant mod $2\pi,$ we find that we need
-$\Phi_{\text{par}}=2\Phi_{\triangle}=\pi\left(2n+1\right),\qquad n\in\mathbb{Z}.$ (59)
+Requiring that this operation leave the flux invariant modulo $2\pi,$ we find that we need
+$2\Phi_{\triangle}=\pi\left(2n+1\right),\qquad n\in\mathbb{Z}.$ (60)
 This is consistent with the observation that the Hofstadter spectrum on the triangular lattice only has bands with opposite energy for these values of $\Phi_{\triangle}$.
 In particular, in this work we consider $\Phi_{\triangle}=\pi/2$.
-We find at each momentum $\bm{k}$ that both the energy and Berry curvature are exactly opposite, with the bands having Chern number ±$\pm 1.$
-In general, the naive particle-hole must be composed with a gauge transformation in order to restore all the hopping elements.
+We find at each momentum $\bm{k}$ that both the energy and Berry curvature are exactly opposite, with the bands having Chern number $\pm 1.$
+For a generic choice of Peierls’ hopping phases consistent with this uniform magnetic flux, the naive particle-hole operation must be composed with a gauge transformation in order to exactly restore all the hopping elements.
 Given the particular form of the hoppings Eq.
 (SI) to Eq.
-(SI) for the two-band model, we obtain the following form of the particle-hole symmetry:
+(SI) for the two-band model, one may easily confirm that the following form of the particle-hole operation leaves the Hamiltonian invariant:
 $\mathcal{P}c^{\dagger}(\bm{r})\mathcal{P}^{-1}=(-1)^{x+y}c^{\mathsf{T}}(\bm{r}),\qquad\mathcal{P}i\mathcal{P}^{-1}=+i,$ (61)
-where $\mathsf{T}$ is the transposition operation on the row vector of electronic annihilation operators.
-SII.3 Mirror composed with time-reversal
-The orbital magnetic field breaks the time-reversal and mirror symmetries enjoyed by the $\Phi_{\triangle}\equiv 0$ (mod $\pi$) Hubbard models on the triangular lattice.
-However, the combination of mirror with time-reversal remains a symmetry and should not be spontaneously broken by the CSL or IQH phases.
-In particular, we focus on $M_{y}\mathcal{T}$ since this symmetry acts “on-site” (i.e., within each ring) when the system is placed on the cylinder.
-Recall that $M_{y}$ should send $\bm{r}\to M_{y}\bm{r}$ (flipping the Cartesian $y$-coordinate) while also mapping $\bm{S}\to(-S_{x},S_{y},-S_{z}).$ On the other hand, time-reversal should flip all components of spin.
-Their combination should therefore map $\bm{S}\to(S_{x},-S_{y},S_{z})$, which is easy to implement due to $\mathcal{T}$ being anti-unitary.
-We claim that the desired symmetry is
-$(M_{y}\mathcal{T})c^{\dagger}(\bm{r})(M_{y}\mathcal{T})^{-1}=c^{\dagger}(M_{y}\bm{r})e^{i(-1)^{x}\pi/4},$ (61)
-where the exponential factor is chosen to restore the phases of the hoppings. We confirm this by explicit calculation:
-$(M_{y}\mathcal{T})\hat{h}(M_{y}\mathcal{T})^{-1}$ $=(M_{y}\mathcal{T})\left(\sum c_{\bm{r}+\bm{\delta}}^{\dagger}t_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}c_{\bm{r}}\right)(M_{y}\mathcal{T})$ (62)
-$=\sum c_{M_{y}(\bm{r}+\bm{\delta})}^{\dagger}e^{i(-1)^{\delta_{x}+x}\pi/4}t^{*}_{\bm{r}+\bm{\delta}\leftarrow\bm{r}}e^{-i(-1)^{x}\pi/4}c_{M_{y}\bm{r}}$ (63)
-$=\sum c_{\bm{r}+\bm{\delta}}^{\dagger}e^{i(-1)^{x}\left((-1)^{\delta_{x}}-1\right)\pi/4}t^{*}_{M_{y}(\bm{r}+\bm{\delta})\leftarrow M_{y}\bm{r}}c_{\bm{r}},$ (64)
-which can be shown to be respected for Eqs. (SI) to (SI) above.
-It is easy to verify that $M_{y}\mathcal{T}$ is not Kramers. However, it can be convenient to define a Kramers symmetry by performing an additional $\pi$ spin rotation of about the $y$ axis:
-$(M_{y}\mathcal{T}^{\prime})c^{\dagger}(\bm{r})(M_{y}\mathcal{T}^{\prime})^{-1}=c^{\dagger}(M_{y}\bm{r})(i\sigma_{y})e^{i(-1)^{x}\pi/4}.$ (65)
-One then has that
-$(M_{y}\mathcal{T})^{2}c^{\dagger}(\bm{r})(M_{y}\mathcal{T})^{-2}$ $=+c^{\dagger}(\bm{r}),$ (66)
-$(M_{y}\mathcal{T}^{\prime})^{2}c^{\dagger}(\bm{r})(M_{y}\mathcal{T}^{\prime})^{-2}$ $=-c^{\dagger}(\bm{r}).$ (67)
-Furthermore, we note that
-$(M_{y}\mathcal{T})\bm{S}(\bm{r})(M_{y}\mathcal{T})^{-1}$ $=\left(S_{x},-S_{y},S_{z}\right)_{M_{y}\bm{r}},$ (68)
-$(M_{y}\mathcal{T}^{\prime})\bm{S}(\bm{r})(M_{y}\mathcal{T}^{\prime})^{-1}$ $=-\bm{S}(M_{y}\bm{r}).$ (69)
-SIII Symmetries on the cylinder
-To implement charge-flux threading, we work in the real-space formalism with periodic boundary conditions and vector potential modified according to Eq. (24). We then get
-$t_{(x,y+1)\leftarrow(x,y)}^{\phi}$ $=t\exp\left(i\pi x\right)\exp\left(i\phi/L_{y}\right)$ (70)
-$t_{(x+1,y)\leftarrow(x,y)}^{\phi}$ $=t$ (71)
-$t_{(x+1,y-1)\leftarrow(x,y)}^{\phi}$ $=t\exp\left(-i\pi(x+1/2)\right)\exp\left(-i\phi/L_{y}\right),$ (72)
-so that we pick up an additional phase of $\phi/L_{y}$ when $y$ increments by $1$ One finds that the symmetries of the system are sensitive not only to the parity of $L_{y}$, but also to the threaded charge and spin flux.
-SIII.1 Magnetic translations
-Figure S2: Density-density connected correlations for each of the four studied circumferences $L_{y}=3$ to $6$, with the Hubbard interaction tuned to the values at which the correlation length is maximal at the largest available bond dimension, namely $U(L_{y}=3)=11.125$, $U(L_{y}=4)=11.0$, $U(L_{y}=5)=11.6$, and $U(L_{y}=6)=11.4$, respectively.
-The even-circumference cylinders are threaded with $\Phi_{\text{ext}}=\pi/2$, as in the main text.
-When $L_{y}$ is even, the Hamiltonian can be restored by composing the bare $T_{x}$ with a large gauge transformation.
-In contrast, by Fig.
-S1(b), we know that the effect of translations on the fluxes through loops cannot be gauged away when $L_{y}$ is odd.
-This manifests in the fact that $T_{x}^{\text{mag}}$ [see Eq.
-(48)] is ill-defined, which we see by imposing periodic boundary conditions and comparing $\bm{r}+L_{y}\bm{a}_{2}$ with $\bm{r}$; note that
-$\exp\left(\frac{i}{2}\bm{b}_{2}\cdot(\bm{r}+L_{y}\bm{a}_{2})\right)=\exp\left(i\pi L_{y}\right)\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right)\neq\exp\left(\frac{i}{2}\bm{b}_{2}\cdot\bm{r}\right).$ (73)
-SIII.2 Particle-hole
-Now, let’s fix $\Phi_{\text{par}}=2\Phi_{\triangle}=+\pi$ and interrogate whether placing the system on a YC cylinder changes this conclusion.
-Focusing on a particular ring $r$ around the cylinder with flux $\Phi_{r},$ we note that
+where $\mathsf{T}$ is the transposition operation on the row vector of electronic annihilation operators (indexed by spin).
+To obtain the form of the particle-hole symmetry presented in the main text, we compose the version in Eq. (61) with a conventional spin rotation:
+$\mathcal{P}c^{\dagger}_{x,y}\mathcal{P}^{-1}=(-1)^{x+y}c^{\mathsf{T}}_{x,y}(i\sigma_{y}),\quad\mathcal{P}i\mathcal{P}^{-1}=+i.$ (62)
+The spin rotation is chosen so that the spin operator
+$S_{\alpha}(\bm{r})=\frac{1}{2}c^{\dagger}(\bm{r})\sigma_{\alpha}c(\bm{r})$ (63)
+is left invariant.
+To see this, we note that the unitarity of $\mathcal{P}$ implies that $\mathcal{P}c(\bm{r})\mathcal{P}^{-1}=(-1)^{x+y}(-i\sigma_{y})c^{\dagger\mathsf{T}}(\bm{r})$.
+This allows us to explicitly compute:
+$\mathcal{P}S_{\alpha}(\bm{r})\mathcal{P}^{-1}$ $=\frac{1}{2}\mathcal{P}c^{\dagger}(\bm{r})\mathcal{P}^{-1}\sigma_{\alpha}\mathcal{P}c(\bm{r})\mathcal{P}^{-1}=\frac{1}{2}c^{\mathsf{T}}(\bm{r})\sigma_{y}\sigma_{\alpha}\sigma_{y}c^{\dagger\mathsf{T}}(\bm{r}).$ (64)
+When $\alpha\in\{x,z\}$ we acquire a minus sign but then obtain another from fermion anti-commutation when undoing the transposes.
+On the other hand, when $\alpha=y$ the Pauli matrices commute but the minus sign from fermion anti-commutation cancels with the fact that $\sigma_{y}$ is an anti-symmetric matrix:
+$\mathcal{P}S_{y}(\bm{r})\mathcal{P}^{-1}=\frac{1}{2}c_{s}(\bm{r})(\sigma_{y})_{ss^{\prime}}c_{s^{\prime}}^{\dagger}(\bm{r})=-\frac{1}{2}c_{s^{\prime}}^{\dagger}(\bm{r})(\sigma_{y})_{ss^{\prime}}c_{s}(\bm{r})=+S_{y}(\bm{r}).$ (65)
+SIII Fate of Symmetries on Cylinder
+SIII.1 Particle-hole
+Let us elucidate when placing the system on a YC-$L_{y}$ cylinder explicitly breaks particle-hole symmetry.
+We may use the naive definition of particle-hole Eq.
+(55) because the fluxes in question are gauge-invariant quantities.
+Focusing on a particular cylinder ring $r$ penetrated with $\Phi_{r}$ flux, note that
 $\tilde{\mathcal{\mathcal{P}}}:\Phi_{r}\mapsto\begin{cases}-\Phi_{r}&L_{y}\text{ even},\\ -\Phi_{r}+\pi&L_{y}\text{ odd}.\end{cases}$ (66)
-Therefore, $\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to a gauge transformation) when
+Therefore, $\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to a gauge transformation) when, for all rings $r$, we have that
 $\Phi_{r}\in\begin{cases}\pi\mathbb{Z}&L_{y}\text{ even},\\ \frac{\pi}{2}+\pi\mathbb{Z}&L_{y}\text{ odd},\end{cases}$ (67)
-for all rings $r.$
-When $L_{y}$ is even, Gauss’s law dictates that the flux $\Phi_{r_{1,2}}$ through two adjancent rings $r_{1,2}$ must be equal — note that this allows $T_{x}$ to be a symmetry (up to a gauge transformation).
-Therefore, possibly after threading $\phi$ external flux through each ring of the cylinder, it is possible to tune all the $\Phi_{r}$ simultaneously (to 00 or $\pi$) so that $\tilde{\mathcal{\mathcal{P}}}$ is respected.
+When $L_{y}$ is even, Gauss’s law dictates that the fluxes $\Phi_{r_{1}}$ and $\Phi_{r_{2}}$ through two adjacent rings $r_{1,2}$ must be equal.
+Note that this allows $T_{x}$ to be a symmetry (up to a gauge transformation).
+Therefore, possibly after threading $\phi$ external flux through each ring of the cylinder, it is possible to tune all the $\Phi_{r}$ simultaneously (to 0 or $\pi$) so that $\tilde{\mathcal{\mathcal{P}}}$ is respected.
 Only in those cases is $T_{x}\tilde{\mathcal{\mathcal{P}}}$ also a symmetry (up to gauge).
@@
 This always rules out $T_{x}$ as a symmetry.
-However, so long as $\Phi_{r_{1}}$ is $\pi/2$ or $-\pi/2,$ both Eqs.
+However, so long as $\Phi_{r_{1}}$ is $\pi/2$ or $\pi/2,$ both Eqs.
 (67) and (68) can be respected, so that $\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to gauge).
-However, odd $L_{y}$ is special in that if we tune $\Phi_{r_{1}}$ to 0 or $\pi,$ then $T_{x}\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to gauge) even though neither $T_{x}$ nor $\tilde{\mathcal{\mathcal{P}}}$ are themselves symmetries.
-Appendix D Observable and Correlation Functions
-SI Heisenberg Operator in terms of Fermions
-If the sites are spinless FermionSites and up and down spins are encoded in adjacent FermionSites, we can compute spin correlations indirectly using the fermion creation and annihilation operators.
-Since the spin correlation operator $\vec{S}_{i}\cdot\vec{S}_{j}$ involves the product of two spin operators, we need to express it in terms of the fermion operators.
-In a system with spinless FermionSites, we can define the spin operators as:
-$S_{i}^{+}=c_{i\uparrow}^{\dagger}c_{i\downarrow},\quad S_{i}^{-}=c_{i\downarrow}^{\dagger}c_{i\uparrow},\quad S_{i}^{z}=\frac{1}{2}(c_{i\uparrow}^{\dagger}c_{i\uparrow}-c_{i\downarrow}^{\dagger}c_{i\downarrow}),$ (77)
-where $c_{i\uparrow}^{\dagger}$ creates an up-spin fermion on site $i$ and $c_{i\downarrow}^{\dagger}$ creates a down-spin fermion on site $i$.
-The spin correlation operator $\vec{S}_{i}\cdot\vec{S}_{j}$ can be written in terms of these fermion operators as:
+However, odd $L_{y}$ is special in that if we tune $\Phi_{r_{1}}$ to 0 or $\pi,$ then $T_{x}\tilde{\mathcal{\mathcal{P}}}$ is a symmetry (up to gauge) even though neither $T_{x}$ nor $\tilde{\mathcal{\mathcal{P}}}$ is individually a symmetry.
+SIII.2 Magnetic translations
+As discussed above, when $L_{y}$ is even, the external magnetic flux through all the plaquettes and loops of the cylinder are consistent with a translation symmetry.
+In particular, this generally requires composing a bare translation with a gauge transformation.
+In contrast, the effect of translations on the fluxes through loops cannot be gauged away when $L_{y}$ is odd (see Fig.
+S1b).
+Mathematically, this manifests in the fact that $T_{x}$ [see Eq.
+(49)] is ill-defined, i.e., not single-valued in the coordinate $y$.
+Thus, any gauge transformation which modifies $T_{x}$ and $T_{y}$ while preserving the property that they anti-commute will also fail to be single-valued.
+Appendix D Entanglement spectrum
+Figure S3: Entanglement data at $L_{y}=6$ with $\Phi_{\text{ext}}=\pi/2$, taken at the cut to the right of the $x=0$ ring.
+(a) Weight of the Schmidt values resolved by spin $S_{z}$ and charge $Q$ quantum numbers (see Appendix Sec.
+D for details) at $U/t=6$.
+(b) Entanglement eigenvalues vs.
+circumferential momentum $k_{y}$; colors (with small artificial horizontal separation) indicate spin quantum number (see legend).
+Left panel: for $Q=0e$, with circled numerals indicating the number of nearly-degenerate $S_{z}=0$ entanglement eigenvalues at each momentum.
+Right panel: likewise for $Q=1e$, with numerals counting the number of nearly-degenerate $S_{z}=1/2$ states.
+(c,d) Same for $U/t=14$.
+Here we study the real-space entanglement spectrum on infinite-length cylinders at a partition into two semi-infinite halves.
+The MPS ansatz allows for the efficient computation of the ground state Schmidt decomposition:
+$|\Psi\rangle=\sum_{\alpha}\lambda_{\alpha}|\Phi_{\alpha L}\rangle|\Phi_{\alpha R}\rangle,$ (69)
+where $|\Phi_{\alpha L/R}\rangle$ are orthonormal bases for the left/right Hilbert spaces.
+The Schmidt values are non-negative and normalized as $\sum_{\alpha}\lambda_{\alpha}^{2}=1,$ with $\lambda_{\alpha}^{2}$ being the probability of measuring subsystem $R$ in the state $|\Phi_{\alpha R}\rangle$ (which collapses $L$ to $|\Phi_{\alpha L}\rangle$).
+Tracing out $R$, one obtains a mixed state for $L$ with von Neumann entanglement entropy
+$S=-\sum_{\alpha}\lambda_{\alpha}^{2}\log\lambda_{\alpha}^{2}.$ (70)
+In Ref.
+[50], Li and Haldane conjectured that the entanglement spectrum at a bipartition $L\cup R$ of a quantum Hall ground state encodes the low-energy spectrum of the half-system with a physical edge at the location of the bipartition.
+Given the reduced density matrix $\rho_{L}$, the entanglement spectrum is defined as the eigenvalues of the non-negative-definite Hermitian matrix $H_{L}$ with $\rho_{L}=e^{-H_{L}}$, namely $\epsilon_{\alpha}=-\log\lambda_{\alpha}^{2}$.
+The Li-Haldane proposal has been further investigated in QH systems [80, 53, 55], demonstrated exactly for free fermions [52], and justified more generally using CFT [54, 56, 57].
+The entanglement eigenstates can be labeled by their charge, spin, and circumferential momentum quantum numbers.
+This enables us to define the Schmidt weight in sector $q=(Q,S_{z},k_{y})$ as the restricted sum $W_{q}=\sum_{\alpha|q}\lambda_{\alpha}^{2}$ over states $\alpha$ with quantum number $q$, so that $\sum_{q}W_{q}=1$.
+In the infinite geometry, relative charges are unambiguous while the absolute charge is subtle [81].
+For concreteness, we use the following convention: we fix the origin $(Q,S_{z})=(0,0)$ so that $Q=0$ has the largest Schmidt weight and the average of $S_{z}$ is zero over the distribution.
+In each fixed $Q$ sector, we further place the lowest-lying entanglement eigenvalue at $k_{y}=0$ for our plots.
+In Fig.
+S3(a,c) we plot the Schmidt weight resolved by $Q$ and $S_{z}$ for the $L_{y}=6$ cylinder ground states at $U/t=6,14$.
+We perform the bipartition to the right of ring $x=0$.
+It is found to be strongly peaked at a single $Q$ for the $U/t=14$ state, but not as a function of $S_{z}$.
+This indicates weak charge fluctuations (i.e., Mott insulator) and relatively large spin fluctuations.
+Fig.
+S3(b), we fix the two most prominent values of $Q$ at $U/t=6$ and plot the ES as a function of $k_{y}$; these spectra carry integer vs.
+half-integer spin representations, consistent with being related by a unit charge (i.e., an electron).
+Fig.
+S3(d) shows the ES for the $U/t=14$ CSL state at $Q=0$, the dominant charge sector.
+An equivalent ES is obtained at every other cut along the cylinder, indicating the lack of translation SSB.
+In Fig.
+S4 we plot analogous data for the $L_{y}=5$ cylinder with $\Phi_{\text{ext}}=0$, which has the glide-particle-hole symmetry.
+Within the IQH phase at $U/t=6$, the distribution of Schmidt weights is reflected as a function of $Q$, but nonetheless equivalent, between bond 0 and 1.
+In contrast, the ES in the $U/t=14$ CSL state differs dramatically at bond 0 vs.
+1, with the lowest-lying states carrying half-integer vs.
+integer spin, indicating SSB of $\mathcal{P}T_{x}$.
+Figure S4: Entanglement data at $L_{y}=5$ with $\Phi_{\text{ext}}=0$, analogous to Fig.
+S3.
+(a) For the cut to the right of ring $x=0$ (i.e., on bond 0), at $U/t=6$.
+Entanglement eigenvalues shown in both the $Q=0e$ (upper right panel) and $Q=+1e$ (lower right) sectors.
+(b) Likewise at bond 1; entanglement spectrum shown for $Q=0e$ and Q=1eQ=-1e sectors.
+(c,d) Similarly for $U/t=14.$
+Appendix E Operators in the hybrid $x\,k_{y}$ representation
+In our analysis of the transition, it is useful to study nearest-neighbour hoppings and their correlations.
+Since the states of interest are translation-invariant, namely around the cylinder circumference, we are motivated to study the ring-averaged correlation functions.
+For each nearest-neighbour lattice displacement
+$\bm{\delta}\in\{\pm\bm{a}_{1},\pm\bm{a}_{2},\pm(\bm{a}_{1}-\bm{a}_{2})\},$ (71)
+let us define
+$d^{\dagger}_{\bm{\delta}}(x)$ $=\frac{1}{L_{y}}\sum_{y}c^{\dagger}(x,y)c((x,y)+\bm{\delta}).$ (72)
+By Eq. (35),
+$d^{\dagger}_{\bm{\delta}}(x)$ $=\frac{1}{L_{y}^{2}}\sum_{y}\sum_{k,k^{\prime}}e^{iky}e^{-ik^{\prime}(y+\delta_{y})}c^{\dagger}(x,k)c(x+\delta_{x},k^{\prime})$ (73)
+$=\frac{1}{L_{y}^{2}}\sum_{y}\sum_{k,k^{\prime}}e^{i(k-k^{\prime})y}e^{-ik^{\prime}\delta_{y}}c^{\dagger}(x,k)c(x+\delta_{x},k^{\prime})$ (74)
+$=\frac{1}{L_{y}}\sum_{k}e^{-ik\delta_{y}}c^{\dagger}(x,k)c(x+\delta_{x},k),$ (75)
+so that
+$d^{\dagger}_{\bm{a}_{1}}(x)$ $=\frac{1}{L_{y}}\sum_{k}c^{\dagger}(x,k)c(x+1,k)$ (76)
+$d^{\dagger}_{-\bm{a}_{1}}(x)$ $=\frac{1}{L_{y}}\sum_{k}c^{\dagger}(x,k)c(x-1,k)$ (77)
+$d^{\dagger}_{\bm{a}_{2}}(x)$ $=\frac{1}{L_{y}}\sum_{k}e^{-ik}c^{\dagger}(x,k)c(x,k)$ (78)
+$d^{\dagger}_{-\bm{a}_{2}}(x)$ $=\frac{1}{L_{y}}\sum_{k}e^{+ik}c^{\dagger}(x,k)c(x,k),$ (79)
+and similarly for $\bm{\delta}=\pm(\bm{a}_{1}-\bm{a}_{2})$.
+SI Staggered Heisenberg Order Parameter
+The following documents the numerical computation of the staggered Heisenberg order parameter in the hybrid $xk_{y}$ fermion representation.
+The spin correlation operator $\vec{S}_{i}\cdot\vec{S}_{j}$ can be expressed in terms of fermionic operators by first noting that
+$S_{i}^{+}=c_{i\uparrow}^{\dagger}c_{i\downarrow},\quad S_{i}^{-}=c_{i\downarrow}^{\dagger}c_{i\uparrow},\quad S_{i}^{z}=\frac{1}{2}(c_{i\uparrow}^{\dagger}c_{i\uparrow}-c_{i\downarrow}^{\dagger}c_{i\downarrow}).$ (80)
+The spin correlation operator $\vec{S}_{i}\cdot\vec{S}_{j}$ can then be written in terms of these fermion operators as:
 $\vec{S}_{i}\cdot\vec{S}_{j}=\frac{1}{2}\left(c_{i\uparrow}^{\dagger}c_{i\downarrow}c_{j\downarrow}^{\dagger}c_{j\uparrow}+c_{i\downarrow}^{\dagger}c_{i\uparrow}c_{j\uparrow}^{\dagger}c_{j\downarrow}\right)+S_{i}^{z}S_{j}^{z}.$ (81)
-SII Staggered Heisenberg Order Parameter
-Here we express the staggered-Heisenberg order parameter in the hybrid $(x,k_{y})$ operator representation, where $x$ indexes cylinder rings.
+Next, we express the staggered-Heisenberg order parameter in the hybrid $(x,k_{y})$ operator representation, where $x$ indexes cylinder rings.
 Because of translation invariance around the cylinder, it is useful to define the $y$-averaged inter-ring NN Heisenberg operator
@@
 $S^{a}(x,y)=c^{\dagger}(x,y)\frac{\sigma^{a}}{2}c(x,y)=\frac{1}{L_{y}}\sum_{k,k^{\prime}}e^{i(k-k^{\prime})y}c^{\dagger}(x,k)\frac{\sigma^{a}}{2}c(x,k^{\prime}).$ (83)
-Since both the IQH and CSL ground states exhibit a two-ring translation invariance (i.e., are invariant under $(T_{x})^{2}$ in addition to $T_{y}$) then we can measure the local order parameter
-$\mathcal{O}^{\delta\text{Heis}}_{0}=\mathcal{O}_{\text{H}}(1)-\mathcal{O}_{\text{H}}(0).$ (81)
 In general, the hybrid representation reads
 $\mathcal{O}_{\text{H}}(x)$ $=\frac{1}{L_{y}}\sum_{y}\sum_{a=x,y,z}S^{a}(x,y)S^{a}(x+1,y)$
-$=\frac{1}{L_{y}^{2}}\sum_{a=x,y,z}\sum_{k,k^{\prime}}\sum_{p,p^{\prime}}\left(\frac{1}{L_{y}}\sum_{y}e^{i(k-k^{\prime}+p-p^{\prime})y}\right)c^{\dagger}(x,k)\frac{\sigma^{a}}{2}c(x,k^{\prime})\cdot c^{\dagger}(x+1,p)\frac{\sigma^{a}}{2}c(x+1,p^{\prime})$
+$=\frac{1}{L_{y}^{2}}\sum_{a=x,y,z}\sum_{k,k^{\prime}}\sum_{p,p^{\prime}}\left(\frac{1}{L_{y}}\sum_{y}e^{i(k-k^{\prime}+p-p^{\prime})y}\right)$
+$\qquad\quad\times c^{\dagger}(x,k)\frac{\sigma^{a}}{2}c(x,k^{\prime})\cdot c^{\dagger}(x+1,p)\frac{\sigma^{a}}{2}c(x+1,p^{\prime})$
 $=\frac{1}{L_{y}^{2}}\sum_{a=x,y,z}\sum_{k,k^{\prime}}\sum_{p,p^{\prime}}c^{\dagger}(x,k)\frac{\sigma^{a}}{2}c(x,k^{\prime})\cdot c^{\dagger}(x+1,p)\frac{\sigma^{a}}{2}c(x+1,k-k^{\prime}+p).$
-Appendix E Correlation Lengths
-Figure S3: Correlation length data for even-circumference geometries.
-Left panel: Illustration of the external magnetic flux penetrating each loop of the infinite cylinder, namely no flux, $\Phi=0$.
-Center panels: Maximum correlation length at $L_{y}=4$ in the $Q=0e$ charge sector [blue] for a range of bond dimensions increasing with increasing darkness; and in the $Q=1e$ charge sector [green].
-Right panel: The same for $L_{y}=6$.
-Figure S4: Correlation length data for even-circumference geometries.
-Left panel: Illustration of the external magnetic flux penetrating each loop of the infinite cylinder, which alternates between ±$\pm\pi/2$, corresponding to threaded flux $\Phi_{\text{ext}}=\pi/2$ in our conventions.
-Center panels: Maximum correlation length at $L_{y}=3$ in the $Q=0e$ charge sector [blue] for a range of bond dimensions increasing with increasing darkness; and in the $Q=1e$ charge sector [green].
-Right panel: The same for $L_{y}=5$.
-Here we briefly comment on additional data concerning the correlation lengths for geometries not displayed in the text.
+Figure S5: Left panel: Sketch of three rings of cylinder at indicated $x$ values (cylinder ring indices), with solid and dashed lines showing along which staggered-Heisenberg order parameter [defined in main text] changes under glide particle-hole symmetry $PT_{x}$ on odd-circumference cylinders.
+Right panel: Magnitude of the staggered Heisenberg order parameter of Eq.
+(84) plotted against $U/t$ for cylinder widths indicated and with $PT_{x}$ preserving and broken phases indicated by the check and cross symbols.
+Consistent with the recovery of this symmetry at the putative 2+1D critical point, the magnitude of the order parameter [bottom left] decreases substantially from $L_{y}=3$ to 5.
+Since both the IQH and CSL ground states exhibit a two-ring translation invariance (i.e., are invariant under $(T_{x})^{2}$ in addition to $T_{y}$) then we can measure the following operator defined on the first two bonds near the origin:
+$\mathcal{O}^{\delta\text{Heis}}_{0}=\mathcal{O}_{\text{H}}(0)-\mathcal{O}_{\text{H}}(1).$ (84)
+In Fig.
+S5, we plot this Heisenberg order parameter on the $L_{y}=3,5$ geometries with $\Phi_{\text{ext}}=0$, which have the glide-particle-hole symmetry.
+This is an alternative to the order parameter computed and displayed in in Fig.
+3(a) of the main text, and defined in Eq.
+(3), which is instead based on the circumferential current operator.
+Both order parameters demonstrate that $\mathcal{P}T_{x}$ symmetry is spontaneously broken in the CSL phase on these two geometries.
+Appendix F Even-circumference supplementary data
+SI Reproducing correlation lengths of Kuhlenkamp et al.
+Figure S6: (a) Plot of the maximum correlation length in the charge-$0e$ [middle panel] and charge-$1e$ [right panel] sectors of the ground state transfer matrix for the YC6 geometry (at the particle-hole-symmetric value $\Phi_{\text{ext}}=0$), with bond dimensions increasing light-to-dark [see legend].
+Left illustration: depiction of the flux through the various cylinder rings, consistent with no externally-threaded magnetic flux.
+(b) Same for YC4, likewise at $\Phi_{\text{ext}}=0$.
+Here we briefly comment on additional data concerning the correlation lengths for geometries not already highlighted in the main text.
 In particular, Fig.
-S3 shows correlation length data for the $L_{y}=4$ and $6$ cylinders which are threaded by no external flux.
-In particular, $L_{y}=6$ reproduces the behavior presented in Kuhlenkamp et al., [11].
-We remark that the correlation lengths are quantitatively smaller than those at $\Phi_{\text{ext}}$ external flux, which are shown in the main text.
-We similarly show data for the $L_{y}=3$ and $5$ cylinders threaded by $\Phi_{\text{ext}}=\pi/2$ external flux.
-Unlike at $\Phi_{\text{ext}}=0$ in the main text, the flux breaks the glide-PH symmetry, and thus we see no diverging or quantitatively large correlation length (as is consistent with no spontaneous symmetry breaking).
-Appendix F Correlation Length Extrapolation
-Figure S5: (a) Extrapolated inverse correlation length in the charge-neutral sector at $L_{y}=4$ and $\Phi_{\text{ext}}=\pi/2$ obtained following the method in Ref.
+S6 shows correlation length data for the $L_{y}=4$ and $6$ cylinders which are not threaded by external flux.
+In particular, $L_{y}=6$ reproduces the behavior presented in Kuhlenkamp et al., [19], pushed here to bond dimension $\chi=32768$.
+We remark that the correlation lengths are smaller than those with $\Phi_{\text{ext}}=\pi/2$ external flux exhibited in the main text, but similarly exhibit a maximum in the correlation length in the $(Q,S_{z})=(0e,0\hbar)$ sector.
+The peak (whose height is still increasing with bond dimension) is located at $U/t=10.6$ at the largest bond dimension $\chi=32768$; the $U$ location of the peak is also still slowly growing with bond dimension.
+On the $L_{y}=4$ cylinder, the correlation length in the $(Q,S_{z})=(0e,0\hbar)$ sector similarly exhibits a finite, local maximum.
+It is peaked at the smaller value $U/t=10.0$.
+There is an uptick in the correlation length at $U/t=11.7$, which is accompanied by an identical value in the $(Q,S_{z})=(0e,\pm\hbar)$ sectors; this feature indicates the softness of a spin-triplet excitation at $L_{y}=4$, and is sub-dominant in comparison at $L_{y}=6$.
+We note that the $Q=1e$ correlation length does not show any feature in the vicinity of this peak or at the uptick.
+SII Flux-threaded system: correlation length extrapolation for main text
+Figure S7: (a) Inverse correlation length in the charge-neutral sector at $L_{y}=4$ and $\Phi_{\text{ext}}=\pi/2$.
+The $y$-intercept of the fitted curve is the correlation length extrapolated to infinite bond dimension; the curve is obtained following the method in Ref.
 [64], as detailed in Sec.
-F, for a range of Hubbard strengths [different colors] and for a sequence of increasing bond dimensions [same-colored points moving sequentially to the left].
-The extrapolated inverse correlation length saturates at a minimum value exceeding $0.15$; the correlation length is equivalently bounded above by $6.6$ ring spacings.
+SII, for a range of Hubbard strengths [different colors], sampled from a sequence of increasing bond dimensions [same-colored points moving sequentially to the left].
+The extrapolated inverse correlation length saturates at a minimum value exceeding $0.15$, i.e., the correlation length is bounded above by $6.6$ ring spacings.
 (b) The same for $L_{y}=5$ and $\Phi_{\text{ext}}=0$.
-In this case, the extrapolated inverse correlation length reaches and descends beyond 0, signaling the need for larger bond dimensions to accurately estimate large correlation lengths.
 To obtain the extrapolated correlation lengths shown in the main text, we employ a recently-developed method for extrapolating correlation lengths in DMRG [64, 82].
-Namely, we compute several of the largest eigenvalues of the transfer matrix $TX_{i}=\lambda_{i}X_{i}$ in the electric charge $Q_{E}=1$ sector for each $k_{y}$, and select those from the same “excitation branch” [52], namely carrying a consistent complex phase.
-These are related to the correlation lengths as
-$\epsilon_{i}=1/\xi_{i}=-\log\left|\lambda_{i}\right|$ (82)
+Namely, we compute several of the largest eigenvalues of the transfer matrix $TX_{i}=\lambda_{i}X_{i}$ in a given charge (i.e., electric charge and spin) sector for each $k_{y}$, and select those from the same “excitation branch” [64], namely carrying a consistent complex phase.
+These are related to the correlation lengths by
+$\epsilon_{i}=1/\xi_{i}=-\log\left|\lambda_{i}\right|,$ (85)
 with $\epsilon_{i}\leq\epsilon_{i+1}$.
-We report $\xi$ in units of cylinder rings, or equivalently $u=a\sqrt{3}/2$ where $a$ is the lattice spacing of the underlying triangular lattice.
+Throughout this work, we report $\xi$ in units of cylinder rings, or equivalently $u=a\sqrt{3}/2$ where $a$ is the lattice spacing of the underlying triangular lattice.
 Correspondingly, $\epsilon$ is reported in units of $u^{-1}$.
-For a system that is gapped in a charge sector $q=(Q,S_{z},k_{y})$, both $\epsilon_{1}^{q},\epsilon_{2}^{q}$ converge to the same infinite value as $\chi\to\infty$, namely the finite inverse correlation length.
+For a system that is gapped in a charge sector $q=(Q,S_{z},k_{y})$, both $\epsilon_{1,q},\epsilon_{2,q}$ converge to the same finite value as $\chi\to\infty$, namely the finite inverse correlation length.
 As a result, the quantity
-$\delta^{q}=\epsilon_{2}^{q}-\epsilon_{1}^{q}$ (83)
+$\delta_{q}=\epsilon_{2,q}-\epsilon_{1,q}$ (86)
 vanishes at $\chi\to\infty$, serving as a good “scaling variable” that measures the deviation from convergence.
-To estimate $\xi(\chi\to\infty)$, we therefore compute $\epsilon(\delta)$ at each available $\epsilon(\chi),\delta(\chi)$ and extrapolate to $\delta=0$ with a polynomial fit, namely
+To estimate $\xi(\chi\to\infty)$, we therefore compute $\epsilon(\delta)$ at each $\chi$ (which gives a data point $(\epsilon(\chi),\delta(\chi))$) and extrapolate to $\delta=0$ by fitting to the form
 $\epsilon(\delta)=a+b\delta^{c}.$ (87)
 This extrapolation is shown in Fig.
-S5 for both $L_{y}=4$ at $\Phi_{\mathrm{ext}}=\pi/2$ and $L_{y}=5$ at $\Phi_{\mathrm{ext}}=0$, yielding the extrapolated values shown in the main text.
-In both cases, the correlation length is computed in the $(Q,S_{z},k_{y})=\bm{0}$, which hosts largest correlation length.
-Appendix G Density-density correlations at all circumferences
-Figure S6: Density-density connected correlations for each of the four studied circumferences $L_{y}=3$ to $6$, with the Hubbard interaction tuned to the values at which the correlation length is maximal at the largest available bond dimension, namely $U(L_{y}=3)=11.125$, $U(L_{y}=4)=11.0$, $U(L_{y}=5)=11.6$, and $U(L_{y}=6)=11.4$, respectively.
-The even-circumference cylinders are threaded with $\Phi_{\text{ext}}=\pi/2$, as in the main text.
-In Fig. S6 we show the ring-averaged density-density connected correlations, defined as $\langle n_{\text{av}}(x)n_{\text{av}}(0)\rangle,$ where
-$n_{\text{av}}(x)$ $=\frac{1}{L_{y}}\sum_{y}c^{\dagger}(x,y)c(x,y)=\frac{1}{L_{y}}\sum_{k}c^{\dagger}(x,k)c(x,k).$ (85)
-Panels (a,c), which correspond to the odd circumferences $L_{y}=3,5$ respectively, exhibit apparently algebraic correlations at long distances.
-On the other hand, panels (b,d) exhibit exponentially-decaying correlations that nonetheless feature a near-algebraic plateau, over a range of $x$ less than their respective correlation lengths.
+S7 for $L_{y}=4$ at $\Phi_{\mathrm{ext}}=\pi/2$, yielding the extrapolated values shown in the main text.
+The correlation length is computed in the $(Q,S_{z},k_{y})=0$ charge sector, which hosts largest correlation length.
+We also perform the extrapolation for $L_{y}=5$ at the glide-particle-hole symmetric flux $\Phi_{\mathrm{ext}}=0$, which instead exhibits a divergence in $\xi_{0}$ at a value of $U$ in the range $U\in(11.4,11.8)$; In the main text we take $U_{c}(L_{y}=6)=11.6$, at which the correlation length is largest at the largest bond dimension $\chi=16384$.
+Appendix G Supplementary odd-circumference analysis
+SI YC3 Correlation Lengths
+Figure S8: (a) Plot of the maximum correlation length in the charge-$0e$, charge-$1e$, and charge-$2e$ sectors of the ground state transfer matrix for the YC3 cylinder geometry, with bond dimensions increasing light-to-dark.
+Left illustration: depiction of the flux through the various cylinder rings, consistent with no externally-threaded magnetic flux, $\Phi_{\text{ext}}=0$.
+(b) Same for the case $\Phi_{\text{ext}}=\pi/2$, for which the Hamiltonian explicitly breaks glide-PH symmetry and therefore does not exhibit SSB criticality.
+In Fig.
+S8, we display the correlation lengths for the YC3 geometry.
+In particular, in panel (a) we provide data for the $\Phi_{\text{ext}}=0$ case, which accompanies Fig.
+3 of the main text.
+Here, the system respects glide-PH symmetry and, like the YC5 system with $\Phi_{\text{ext}}=0$ studied in the main text, exhibits a symmetry-unbroken small-$U$ phase and symmetry-broken larger-$U$ phase.
+The transition between then is evidently continuous, featuring a diverging correlation length which reaches almost 600 ring spacings at the maximum bond dimension $\chi=16384$ considered.
+In Fig.
+S8(b), in contrast, threading $\Phi_{\text{ext}}$ explicitly breaks the glide-PH symmetry, which reduces the transition to a crossover with a small and wide peak in the correlation length.
+SII Relation to ionic Hubbard chain
+Here, we make more precise the relationship between the “ionic Hubbard” chain and our model on those YC-$L_{y}$ cylinders (with odd $L_{y}$ and $\Phi_{\text{ext}}=0$) which possess glide-PH symmetry.
+In particular, the former model can be thought of as a truly 1+1D version of the latter, i.e., a bipartite chain as opposed to a quasi-(1+1)D system with the hopping-connectivity of a cylinder.
+We first compare them at the level of symmetries.
+The ionic Hubbard Hamiltonian is defined as [83]:
+$H=-t\sum_{i,\sigma}(c^{\dagger}_{i,\sigma}c_{i+1,\sigma}+h.c.)+U\sum_{i}n_{i,\uparrow}n_{i,\downarrow}+\frac{\Delta}{2}\sum_{i,\sigma}(-1)^{i}n_{i,\sigma}.$ (88)
+The Hamiltonian is invariant under translation by two sites, with $\Delta$ being the potential energy difference between the two sublattice sites in the unit cell.
+Moreover, it is invariant under site-centered inversion (but not bond-centered inversion), time-reversal, and a “duality” symmetry that is the combination of particle-hole with a single-site translation [73].
+These symmetries of the ionic Hubbard chain correspond to the following symmetries of our model: $(T_{x})^{2},$ site-centered $C_{2z}$, $M_{y}\mathcal{T}$, and glide-PH $\mathcal{P}T_{x}$.
+Remarkably, the two models are also similar at the level of their ground states.
+The small-$U$ ground state phase of the ionic Hubbard chain is a fully-symmetric band insulator.
+Is it widely believed that the phase which appears next upon increasing $U$ is a “spontaneously-dimerized” phase which spontaneously breaks the duality and site-centered inversion symmetries [73].
+Moreover, as mentioned in the main text, it is believed that these two states are separated by a 2D Ising critical point.
+This Ising picture has been substantiated both by weak-coupling bosonization arguments [70, 71] and an effective strong-coupling effective spin model formulation [72].
+However, at least one careful numerical study of the transition using finite DMRG finds features inconsistent with Ising criticality [84], though the numerical observations are still consistent with a gap to charged excitations and a gapless neutral sector [85].
+These two phases should be identified with the odd-circumference (symmetry-respecting) IQH and (symmetry-breaking) CSL phases, respectively.
+Appendix H Density-density correlations at all circumferences
+In Fig.
+S9 we show various connected correlation functions for the even-circumference cylinders studied in Fig.
+3 of the main text, namely $L_{y}=4,6$ with $\Phi_{\text{ext}}=\pi/2$.
+We plot the current-current (Fig.
+S9a) and density-density (Fig.
+S9b) connected correlation functions, which extend far beyond the spin-spin (Fig.
+S9c) and electron-hole (Fig.
+S9d) correlators (defined in main text).
+The first two extend to much longer range than the latter, suggesting both the existence of a spin gap across the transition and the prominent of spin-singlet charge-neutral fluctuations at criticality.
+While the even-$L_{y}$ systems always exhibit a crossover from IQH to CSL, their behavior is qualitatively consistent with the odd-$L_{y}$ cylinders with $\mathcal{P}T_{x}$ symmetry: the spin-spin and electron-hole correlations are also rapidly exponentially decaying, and the current-current correlator is likewise dominant.
+Figure S9: Connected correlation functions for the $L_{y}=4,6$ systems with $\Phi_{\text{ext}}=\pi/2$.
+The Hubbard interaction is tuned to the value where the correlation lengths are maximal, namely $U(L_{y}=4)=11.0$ and $U(L_{y}=6)=11.4$.
+(a) Current-current correlator (defined in main text).
+(b) Correlator of ring-averaged charge density.
+(c) Correlator of ring-averaged spin density density.
+(d) Electron-hole correlator (defined in main text).
+The $L_{y}=6$ correlations are reduced by $\times 10$ for ease of viewing.